### v1

# HAFixAgent: History-Aware Automated Program Repair Agent

Abstract.
Automated program repair (APR) has recently shifted toward large language models and agent-based systems, yet most systems rely on local snapshot context, overlooking repository history. Prior work shows that repository history helps repair single-line bugs, since the last commit touching the buggy line is often the bug-introducing one. In this paper, we investigate whether repository history can also improve agentic APR systems at scale, especially for complex multi-hunk bugs. We present HAFixAgent, a History-Aware Bug-Fixing Agent that injects blame-derived repository heuristics into its repair loop. A preliminary study of all 854 real-world bugs from Defects4J motivates our design, showing that bug-relevant history is both widely available and highly concentrated. Empirical comparison of HAFixAgent with two state-of-the-art baselines shows: (1) Effectiveness: HAFixAgent significantly improves over the agent-based baseline (by 212.3%) and the multi-hunk baseline (by 29.9%). (2) Efficiency: history does not significantly increase agent steps and keeps token costs comparable, with notably lower median costs for complex multi-file-multi-hunk bugs. (3) Practicality: combining different historical heuristics repairs more bugs, offering a clear cost-benefit trade-off. HAFixAgent offers a practical recipe for history-aware agentic APR: ground the agent in version control history, prioritize diff-based historical context, and integrate complementary heuristics when needed.
Automated Program Repair, Agentic Software Engineering, Large Language Model, Mining Software Repository, Software Engineering Agent, AI Agent, Bug Fixing
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software testing and debuggingccs: Computing methodologies Artificial intelligence
1. Introduction
Software bugs are an inevitable and costly aspect of software development, leading to system failures, security vulnerabilities, and degraded user experience (BOULDER, 2013; Zhang et al., 2024a; Weiß et al., 2007). Since manually identifying and fixing these bugs consumes significant developer time and resources, automated program repair (APR) has been an active and promising research direction for over a decade, aiming to mitigate this burden by automatically generating patches for detected bugs (Le Goues et al., 2019). APR has evolved from heuristic-based (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018), constraint-based (Le et al., 2017; Mechtaev et al., 2016; Long and Rinard, 2015), and template-based techniques (Liu et al., 2019b; Martinez and Monperrus, 2016; Liu et al., 2019a) to learning-based methods (Chen et al., 2021a; Jiang et al., 2021; Li et al., 2020; Lutellier et al., 2020; Zhu et al., 2021, 2023; Ye and Monperrus, 2024) and, most recently, large language model-based approaches (Zhang et al., 2024b; Jiang et al., 2023; Hossain et al., 2024; Xia et al., 2023b; Xia and Zhang, 2024; Xia et al., 2023a; Fan et al., 2023; Yin et al., 2024; Huang et al., 2025; Li et al., 2025a; Shi et al., 2025) and agent-based approaches (Yang et al., 2024; Bouzenia et al., 2025; Mu et al., 2025; Meng et al., 2024; Lee et al., 2024; Zhang et al., 2024c; Chen et al., 2025; Wang et al., 2025). However, despite substantial progress, reliably repairing complex bugs remains a persistent challenge.
Particularly difficult are bugs that span multiple lines and locations (e.g., multi-hunks), which require coordinated, potentially non-local edits across different parts of the codebase. Recent characterizations, such as the work by Nashid et al. (2025), highlight that these bugs are not only common but also significantly harder for current LLM-based approaches compared to simpler single-hunk fixes. Overcoming this complexity barrier is crucial for APR approaches to be broadly applicable in real-world software development.
To address such complex bugs, recent advances in APR have pursued two promising yet separate directions. First, agent-based approaches have emerged as the state-of-the-art (Bouzenia et al., 2025; Yang et al., 2024; Wang et al., 2025). Agent-based approaches like RepairAgent (Bouzenia et al., 2025) leverage LLMs to plan, execute actions via software tools (e.g., code search, editing, testing), and react to iterative feedback, enabling them to tackle more complex reasoning compared to simple prompting. However, these agents typically operate only on the current code snapshot and immediate test results, lacking deeper historical context. Second, drawing from decades of mining software repositories research (Sliwerski et al., 2005; Hassan, 2006; Adams et al., 2010), recent APR techniques started to recognize the value of historical code context. Our prior work (Shi et al., 2025) showed that injecting historical context derived from git blame [^1] https://git-scm.com/docs/git-blame commits (the last change touching a buggy line) can improve LLM repair performance for single-line bugs using simple prompting. However, this approach lacked the sophisticated reasoning and tool interaction capabilities of agents, limiting its potential for complex, multi-location repairs.
We hypothesize that combining the iterative, tool-using capabilities of agentic workflows with the rich, contextual clues found in repository history can significantly enhance APR, particularly for complex multi-hunk bugs. History provides information on relevant past changes, developer rationale, and co-evolution patterns, while the agent leverages tools to explore, validate, and integrate this historical context dynamically during the repair process. Yet, it is unclear whether this approach scales. Complex multi-hunk bugs could have many different blame commits, potentially overwhelming the agent with noisy and irrelevant signals rather than providing a focused context.
We design HAFixAgent, a history-aware agent for APR that integrates history heuristics (inspired by our prior HAFix work (Shi et al., 2025)) to enable the agent to leverage historical context alongside bash command tools during its autonomous execution. To evaluate the effectiveness of HAFixAgent, we apply it to all 854 bugs in the Defects4J (Just et al., 2014) dataset, a widely used benchmark for evaluating APR techniques. In this study, we address the following research questions (RQs):
  • RQ0. To What Extent Is Historical Information Available in Real-World Bugs? Before integrating repository history, we have to first establish if relevant historical information is broadly available and scalable for use in APR. Specifically, we analyze the Defects4J dataset to measure the availability and concentration of historical information for bugs. Our findings confirm that 71.1% of bugs have available historical information and most map to a single unique blame commit. This result directly supports the feasibility of using a simple, single-commit heuristic for HAFixAgent, even for complex bugs.
  • RQ1. How Effective Is HAFixAgent at Fixing Real-World Bugs? We empirically compare HAFixAgent to recent APR systems and also quantify the extent to which any improvements are attributable to the introduction of historical context as opposed to the adoption of an agent architecture. For this, we evaluate HAFixAgent’s effectiveness against two state-of-the-art baselines on the Defects4J dataset: one for common bugs (Bouzenia et al., 2025) and another specializing in multi-hunk repairs (Nashid et al., 2025). The results show that HAFixAgent repairs 212.3% more bugs than the baseline for common bugs and outperforms the multi-hunk baseline by 29.9%.
  • RQ2. What Are the Cost and Efficiency of HAFixAgent? Historical grounding should not inflate agent loops or token usage. We track token costs of HAFixAgent under different configurations and compare the costs. We found that integrating history into the agent repair loop is cost-efficient overall, without notable overhead.
The main contributions of our paper are as follows:
  • The design and implementation of HAFixAgent, a novel history-aware agentic approach for APR that enriches its decision-making loop with deep repository context.
  • A practical and automated method for constructing historical context, featuring three distinct blame-derived historical heuristics and a robust fallback strategy designed for any agentic APR system.
  • A large-scale empirical evaluation on all 854 Defects4J bugs, providing strong evidence that history is (1) broadly available (71.1% blameable) and highly concentrated (with 70.7% of bugs mapping to a single commit), and (2) significantly improves agent effectiveness and efficiency.
  • An open-source and replicable artifact of HAFixAgent (Shi, 2025), including its agent design, historical data extractor, and full evaluation framework, to foster replication and future research.
Paper organization.
Section 2 provides background information. Section 3 presents the preliminary study on history availability and distribution. Section 4 describes our architecture design of HAFixAgent. Section 5 presents our case study design. Section 6 presents the results. Section 7 discusses the possible future work. Section 8 outlines threats to the validity of our study. Section 9 reviews related work. Section 10 concludes this paper.
2. Background
In this section, we introduce the background information for our study.
2.1. Mining Repository History for APR
A persistent challenge in LLM-based APR is providing the model with relevant, high-quality context (Prenner and Robbes, 2024; Sintaha et al., 2023). While many systems focus on the current (buggy) code snapshot, the repository’s version control history offers a rich, longitudinal source of context. Decades of Mining Software Repositories (MSR) research have established that historical data is a powerful asset. Hassan (2006) demonstrated the broad utility of mining repositories to assist developers. More specifically, Sliwerski et al. (2005) showed that analyzing the change-inducing commit can help locate faults. Adams et al. (2010) used history to identify co-evolving files or crosscutting concerns, providing a map of code dependencies not visible in a static snapshot.
Building on these MSR insights, researchers have long attempted to create APR systems that directly leverage history. For example, Le et al. (2016) pioneered this by creating a repair model based on patterns mined from previous, successful human-generated patches. This idea of using bug-to-fix patterns became a staple in pre-LLM repair techniques. Similarly, Kamei et al. (2013) reviewed the state of using historical data for the related task of fixing build failures, confirming the value of this approach.
Even in the modern LLM era, this principle of leveraging history-as-experience is re-emerging to enhance agentic repair. For example, EXPEREPAIR (Mu et al., 2025) designs a dual-memory system that augments an LLM’s context by leveraging historical repair experiences from previously resolved issues from the same repository. Likewise, SWE-Exp (Chen et al., 2025) proposes an experience-driven framework where the agent learns from the past issue-resolution trajectories (i.e., history) to improve its current issue repair.
These agent-based approaches show the value of history but rely on similar bugs or issues occurring before in the code repository or past trajectories. A more direct and practical heuristic is available through tools like git blame. For any given file, git blame annotates each line with the specific commit that last modified it. The commit identified by this command serves as a powerful heuristic. This choice aligns with SZZ’s core assumption: given a bug-fixing change, SZZ blames the lines modified or deleted by the fix in the pre-fix revision and treats the last change to those lines as a likely bug-introducing commit (Sliwerski et al., 2005; Kim et al., 2006). Analyzing the patch (diff) and commit message from this single, relevant change can provide critical clues about the code’s recent evolution and developer intent. Our prior work, HAFix (Shi et al., 2025), validated this exact approach, demonstrating that injecting this blame-derived context significantly improves repair for single-line bugs. This paper builds directly on that finding: we aim to explore the hypothesis that integrating this same focused historical context into a modern agentic workflow (Section 2.2) can scale these benefits to handle more complex, multi-location repairs.
2.2. Agentic Software Engineering
The emergence of powerful large language models (LLMs), such as GPT-5 (OpenAI, 2025a) and DeepSeek-V3 (DeepSeek-AI et al., 2024), has catalyzed a paradigm shift in software engineering, moving from static, one-shot code generation to dynamic, autonomous agent-based systems. An LLM-based agent is a system that couples the core reasoning and planning capabilities of an LLM with essential components like memory (to maintain state) and a set of tools (to interact with an environment) (Wang et al., 2024). Unlike a simple prompt-and-response model, an agent operates within an iterative execution loop, often conceptualized as a Reason-Act (ReAct) cycle (Yao et al., 2023). This loop allows the agent to formulate a plan, execute an action, observe the outcome, and then autonomously react to that feedback to progress toward a complex, high-level goal.
This iterative, stateful, and interactive approach is the foundation of Agentic Software Engineering (ASE), an emerging field that marks a paradigm shift for software engineering in the era of foundation models (Hassan et al., 2024). The vision of ASE is to create autonomous ”AI Teammates” or ”AI Software Developers” that can operate as generalist partners, capable of handling complex, end-to-end software development tasks (Hassan et al., 2025; Li et al., 2025b). The software engineering domain is a particularly fertile ground for agents because developer tasks are goal-oriented (e.g., ”fix this bug,” ”implement this feature”) and require interaction with a rich, well-defined tool ecosystem, including file systems, build tools, compilers, test runners, and version control systems (Roychoudhury, 2025).
Within the broad field of ASE, a primary and highly active research area has been software debugging and repair, specifically for resolving bugs and GitHub issues. This has led to a new generation of agentic APR systems. For example, SWE-agent (Yang et al., 2024) introduced an Agent-Computer Interface (ACI) that equips an agent with simple, developer-centric tools (such as file navigation, editing, and searching) to autonomously resolve real-world GitHub issues. Similarly, the OpenHands (Wang et al., 2025) seeks to build generalist agents that can execute complex tasks by reasoning over a plan and invoking tools. In the specific APR domain, RepairAgent (Bouzenia et al., 2025) was a pioneering work that provided an LLM with a dedicated toolset to autonomously read and extract code, search code, generate patches, and react to test feedback. Around the same time, systems like AutoCodeRover (Zhang et al., 2024c) further enhanced agents by integrating context from program analysis, such as API calls, to improve repair performance.
These systems represent the state-of-the-art in agent-based program repair. However, a common limitation is that the context these agents leverage is typically limited to the current, static snapshot of the codebase and the immediate feedback from the test suite. They generally lack awareness of the codebase’s historical and evolutionary context: how and why files have changed over time. HAFixAgent, the APR system proposed in this paper, is designed to bridge this exact gap by enriching this agentic loop with actionable, historical context derived from version control history.
3. Preliminary Study (RQ0): To What Extent Is Historical Information Available in Real-World Bugs?
In this section, we provide the motivation, approach, and results for the preliminary study.
3.1. Motivation
Providing LLM agents with informative context is essential for effective repair (Prenner and Robbes, 2024). While most context engineering approaches for automated program repair (APR) focus on knowledge obtained from a given snapshot of a code base (Prenner and Robbes, 2024; Haque et al., 2025; Fakhoury et al., 2024; Chen et al., 2024; Xia et al., 2023b; Li et al., 2025a), only recently APR techniques were proposed that leverage code history (Shi et al., 2025; Ehsani et al., 2025). Inspired by the domain of mining software repositories (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013), our prior work on single-line APR leverages context related to the last commit that touched the bug location, as the latter commit commonly is assumed to be the bug-introducing commit and hence has valuable knowledge about the introduced bug. Finding this commit for single-line bugs is straightforward (as obtained using the git blame command), as the latter only comprise one buggy line.
Unfortunately, more complex bugs touching one hunk, multiple hunks in one file or multiple hunks across files potentially would require combining blame commit information of many different lines, potentially drowning the latter information within an LLM’s token window, and hence reducing the performance of historical context for APR. In such a case, an APR technique would need to aggregate and rank multiple blame history sources, increasing prompt size, latency, and the risk of distraction. Furthermore, for a given buggy snapshot of the code base, blame commits do not exist for all buggy lines. For instance, when fault localization identifies the fault location as an insertion point (where code is missing) rather than an existing line, git blame cannot be applied, as it only annotates existing code. Notably, this process focuses solely on the buggy lines identified by fault localization, which does not depend on the actual bug fix.
This preliminary RQ studies how much of a problem the scale of blame history is for single-hunk, multi-hunk and multi-file bugs: (i) how often does a blame commit exist for buggy lines? and (ii) across how many unique commits is the historical blame information of these types of bugs spread? The results will provide clear empirical insights about the effort required to adapt history-based context to more complex types of bugs.
3.2. Approach
3.2.1. Dataset
To study the availability and distribution of historical information in real-world bugs, we choose the latest version of the Defects4J dataset (v3.0.1) (Just et al., 2014), a standard and widely-used benchmark for evaluating recent APR techniques (Xia and Zhang, 2022; Jiang et al., 2023; Hossain et al., 2024; Lutellier et al., 2020; Jiang et al., 2021; Ye and Monperrus, 2024; Bouzenia et al., 2025). It contains 854 Java bugs alongside test cases for each bug collected from 17 different open-source projects, spanning various domains such as data visualization (Chart), compiler construction (Closure), and date/time utilities (Time). Analyzing the entire dataset allows us to evaluate the general history distribution on different projects and bugs, without restricting to small or specific bug categories.
Consistent with common practices in the APR field (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we assume perfect fault localization. Specifically, we use the developer-written patch provided by Defects4J to identify the buggy lines (i.e., the lines in the buggy snapshot that were modified or deleted in the fix). Importantly, our approach does not rely on the actual bug fixes themselves, but solely on the locations of the buggy lines. These identified lines are the only lines on which git blame is subsequently computed. We then run git blame on these lines to obtain the most recent commits modifying those lines. Finally, we de-duplicate the commit hashes obtained for each bug, count each bug’s unique blame commits, and compute the distribution by bug category.
3.2.2. Bug Category Terminology
To categorize bugs based on the structure of their bug fix commits, we first define a hunk as a single, contiguous block of edits (i.e., additions, deletions, or modifications) applied to a specific region in the source code, similar to earlier work (Nashid et al., 2025). A patch file can contain one or more such hunks. Furthermore, following prior work on bug categorization (Nashid et al., 2025; Bouzenia et al., 2025; Jiang et al., 2023; Xia and Zhang, 2024), we classify all 854 bugs into four mutually exclusive categories based on the number and location of these hunks:
  • Single-Line (SL): The bug fix commit edits exactly one line, contained within a single hunk in a single file.
  • Single-Hunk (SH): The bug fix commit edits one contiguous hunk in one file, and consists of two or more line changes.
  • Single-File-Multi-Hunk (SFMH): The bug fix commit contains two or more distinct hunks, but all hunks are confined to a single file.
  • Multi-File-Multi-Hunk (MFMH): The bug fix commit contains hunks that span two or more different files.
We also categorize bugs by blame availability:
  • Blameable: At least one edited line in the bug fix commit is a deletion or modification relative to the buggy code snapshot, so git blame can return the most recent commit touching such lines.
  • Blameless: The bug fix commit only inserts new lines or files, with no deletions or modifications, so no line maps to a prior commit and hence git blame is not applicable.
Unlike prior work that either nests SL within SH (Jiang et al., 2023) or collapses SFMH and MFMH into one multi-hunk class (Nashid et al., 2025), we keep all four categories separate for finer granularity.
Figure 1. Distribution of the number of unique blame commits across different bug categories on Defects4J.
Table 1. Blame availability by bug category in Defects4J, assuming perfect fault localization. Percentages in the Blameable or Blameless columns are within-category; Total shows the share relative to the full dataset.
Category Blameable Blameless Total
SL 140 83.8 27 16.2 167 19.6
SH 70 52.2 64 47.8 134 15.7
SFMH 294 69.2 131 30.8 425 49.8
MFMH 103 80.5 25 19.5 128 15.0
Total 607 71.1 247 28.9 854 100.0
3.3. Results
Historical blame commit information is broadly available, even for complex bugs. As shown in Table 1, the majority (71.1%) of bugs in Defects4J are blameable, i.e., have at least one blame commit across all buggy lines. Notably, complex multi-hunk bugs (SFMH and MFMH), which account for 64.8% of the entire dataset, maintain high history availability: 69.2% of SFMH bugs and 80.5% of MFMH bugs are blameable. Meanwhile, SL has 140 of 167 blameable (83.8%) and SH has 70 of 134 (52.2%). SFMH is the largest subset at 49.8% of all bugs.
These numbers indicate that historical blame data is commonly available across categories, and hence can be used to ground the reasoning of an APR technique. That said, for the almost thirty percent for which no blame commit can be determined, a workaround would need to be provided.
All but three of the blameable Defects4J bugs have exactly one (!) unique blame commit. Figure 1 shows the distribution of unique blame commits across all bugs. Contrary to expectations, 70.7% (604 out of 854) of all bugs map to exactly one unique blame commit To validate this finding, we manually inspected a sample of multi-hunk cases and confirmed that their blameable lines did indeed trace back to a single, common commit. While one would expect multi-hunk bugs to require at least one separate blame commit per hunk, it turns out that the (blame) history of buggy hunk lines overlaps almost exactly. Those blame commits could still change other lines in addition to the buggy lines (potentially diluting blame context for APR), but the need to add context for only one blame commit suggests that adapting this heuristic to more complex bug categories could be feasible!
In contrast, 28.9% (247 out of 854) have no blame commits (the add-only, blameless cases), and only 0.4% (3 out of 854) of the bugs have two or more git blame commits. The same pattern holds within each category, for example, SL has 140 bugs with one blame commit and 27 with none, and MFMH has 102 bugs with one blame commit and 25 with none.
Summary for Preliminary Study (RQ0): (1) Blame commits can be identified for 71.1% of bugs (when assuming perfect fault localization). (2) Even for complex bug types, blame history typically is concentrated into one commit, as 70.7% of Defects4J bugs have exactly one unique blame commit.
Figure 2. HAFixAgent architecture and workflow.
4. HAFixAgent
Based on the findings of the preliminary study, we propose a new APR approach called HAFixAgent that exhibits the following three design decisions:
  1. (1)
    a minimal agent-based architecture designed to iteratively handle complex, multi-step reasoning and actions (e.g., code exploration, patch validation) required for bugs that go beyond simple, single-line fixes.
  2. (2)
    the agent has access to a git blame tool, inspired by the findings in Section 3, and leverages the single and unique blame commit as the primary source of historical context.
  3. (3)
    the agent includes a fallback mechanism to obtain the relevant history for blameless bugs where no buggy line has an associated blame commit.
Unlike agent-based APR systems such as RepairAgent (Bouzenia et al., 2025), which employ pre-defined Python tool APIs, HAFixAgent is intentionally minimal and history-aware. It treats repository history from previous snapshots as an important source of context that augments the current code snapshot, and it operates within a compact, guarded loop over a small set of standard bash tools (e.g., grep, sed, find). This design choice is critical: it minimizes ambiguity from complex and high-level tool interactions, allowing us to more directly attribute performance gains to the quality of the historical context rather than to sophisticated tool exploration. Additionally, the design is consistent with recent commercial coding assistants such as Claude Code (Anthropic, 2025) and OpenAI Codex (OpenAI, 2025b).
Figure 2 overviews the architecture and its three primary modules. The workflow proceeds as follows:
  • Context Builder: This module prepares the full context for the agent. It assembles the Metadata Context 1 and uses its History Extractor to retrieve 2 and format 3 the Historical Context. All context is then organized into the User Prompt and System Prompt 4.
  • Agent Execution Loop: This module controls the iterative repair. It starts with the LLM query 5, which generates an action 6. This action is sent to the Tools module 7. The loop then receives the raw feedback 8 from the tools, tracks the test result 9, and checks termination guards 10 to decide whether to continue. If so, the feedback is routed to the next LLM query $1$, repeating the cycle.
  • Tools: This module provides an isolated sandbox environment for action execution. It receives an action (e.g., search, edit, test) from the execution loop 7, then returns the resulting feedback (e.g., stdout/stderr) to the loop 8.
4.1. Context Builder
4.1.1. Metadata Context
The user prompt contains three items: the human-written bug report, the initial set of failing tests, and the fault localization context, which specifies the buggy files and lines. These metadata are used only as non-history context and are rendered in step 1.
4.1.2. History Extractor
The History Extractor module (step 2 in Figure 2) is responsible for finding a relevant blame commit for every bug. We use the buggy lines identified through the fault localization metadata to obtain the blame commit, since the historical context in that blame commit may reveal the root causes of the bug (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013). The module handles blameable and blameless bugs differently:
  • Blameable bugs. For bugs involving modified or deleted lines, we execute git blame on all identified lines and collect the union of all unique commit hashes. This process is deterministic rather than a random selection of one line. As demonstrated in our preliminary study (Section 3), this union is nearly always a single commit: 70.7% of all bugs (except three blameable bugs) map to exactly one unique blame commit.
  • Blameless bugs (fallback strategy). To enable these blameless bugs (i.e., add-only fixes) to also benefit from historical context, we design a fallback strategy. Specifically, we employ a fallback strategy to blame the nearest executable code line (e.g., an executable line rather than a comment or symbol) within the five lines preceding the insertion. This strategy is based on the assumption that new code is often added near existing, related logic, making the most recent commit to that local area a relevant source of historical context.
4.1.3. Historical Context
Once the blame commits are determined by the History Extractor, step 3 then mines relevant historical context from the identified blame commits by leveraging one of the three top-performing historical heuristics from our HAFix approach (Shi et al., 2025) (which focused on SL bugs only):
  • All functions’ names from the blame commit (fn_all): includes the names of all functions from all co-changed files in the blame commit. This setting captures a broad structural view of the co-evolved functions within the same history commit.
  • Historical function before and after (fn_pair): which contains the before and after snapshot of the function code body that includes the buggy line. This historical context helps the agent understand how the buggy function evolved and provides possible clues about the root cause of the bug.
  • File-level diff (fl_diff): includes the diff patch obtained from the git diff command in the blame commit. This configuration exposes the exact code modification across files, allowing the agent to reason about fine-grained edits that may have caused the bug.
In step 4, HAFixAgent integrates the history context, which was extracted in the previous step 3 using one of the three heuristics from the blamed commit. The history is injected directly into the User Prompt (as shown in Figure 2), with truncation safeguards (e.g., truncating an exceptionally large file diff) applied to ensure that the total prompt fits within the LLM’s context window. This process leads to our four context configurations: the agent operates either with the non-history context only (from Section 4.1.1) or with the non-history context augmented by one of the three history heuristics (fn_all, fn_pair, or fl_diff).
4.2. Agent Execution Loop
HAFixAgent’s execution loop follows an observe-act-feedback paradigm, inspired by frameworks like ReAct (Yao et al., 2023) and those used in recent software engineering agents (Yang et al., 2024). The loop proceeds with the following steps: 5 the LLM receives the current state (including the full conversation history of prior actions and observations); 6 the agent parses the model output to extract a tool invocation (i.e., a bash command, which the system prompt constrains the model always to provide) as the next action; 7 the command runs in a containerized tool sandbox that exposes repair utilities and dataset tools (e.g., search, edit and compile); 8 stdout and stderr are collected from the output of command execution as feedback; 9 compilation and test outcomes are captured and parsed; 10 the loop terminates on success or when a termination guard (e.g., step limit, cost limit, or timeout) is reached; otherwise, $1$ the observation is appended to the conversation history and fed back to the LLM for the next reasoning step.
4.3. Tools
Table 2. Bash tools used by HAFixAgent.
Tool Description Usage Example
Repair Operations
grep Search for patterns in files; Quick file overview of classes, interfaces, and method signatures grep -n "class|interface|public.*(" file.java
sed -n Read specific line ranges from files; extract targeted context around fault locations progressively sed -n ’45,65p’ BuggyClass.java (reads lines 45-65)
sed -i In-place file editing; make precise, targeted fixes to source code sed -i ’s/oldCode/newCode/’ file.java
find Locate files in the repository by name or pattern find . -name "*Test.java"
head / tail Reconstruct files for complex multi-line edits; extract beginning/end of files for rebuilding head -n 50 file.java > temp && cat fix.txt >> temp
Project Commands
compile Compile the project to verify setup and check for compilation errors after edits compile
test Run relevant/failing tests; verify bug fix effectiveness test -r
General Commands
echo Signal task completion when all tests pass echo COMPLETE_REPAIR_SIGNAL
&& / || Chain multiple commands together; execute sequentially or conditionally sed -n ’10,20p’ A.java && sed -n ’30,40p’ B.java
Table 2 lists the bash tools that HAFixAgent may invoke during the execution loop. The agent emits exactly one bash command per step, with optional short chaining via && or ||. These actions are executed between steps 7 and 8 in the loop, and the same toolset is embedded in the system prompt (Appendix A).
HAFixAgent intentionally exposes a simple but powerful set of bash tools together with project commands like compile and test. A key design choice is that the historical context is not an interactive tool, but is instead modularized and prepared by the History Extractor and injected directly into the agent’s context before the execution loop begins. This decouples history retrieval from agent actions, which allows the agent’s toolset to remain simple, modular, scalable and focused on repair-related operations (e.g., read, edit and test).
The tools in Table 2 fall into three groups. This specific set of tools was chosen to be minimal yet complete, providing the fundamental capabilities (file operations, project compilation, and testing) required for the iterative repair process while relying on universal, standard Linux utilities. (i) File operations: grep gives a quick structural overview, sed -n reads targeted windows around fault locations with progressive expansion, sed -i applies precise edits, head/tail supports multiline reconstruction by splicing, and find locates candidate files. (ii) Project commands: compile command checks build health after edits; test command runs only the relevant tests to provide fast feedback in the execution loop. (iii) General commands: a completion sentinel (echo COMPLETE_REPAIR_SIGNAL) signals success, and simple chaining helps combine a small read or edit with an immediate compile or test.
In practice, the loop alternates between these actions: inspect and read around the localized lines, apply a minimal edit, compile, run relevant tests, and iterate until all tests pass. Termination occurs in two ways: (1) the agent, after running a successful test command, self-terminates by emitting the echo COMPLETE_REPAIR_SIGNAL, or (2) the agent execution loop detects that a guard (step, cost, or timeout) has been reached and halts the process.
5. Case Study Design
In this section, we describe the case study design for evaluating HAFixAgent.
5.1. Prompt Context Preparation
Consistent with the preliminary study, we employ the entire 854 bugs of Defects4J to prepare the data needed for evaluation.
Metadata collection.
To prepare the non-history context as defined in Section 4.1.1, we need bug metadata such as bug reports, failing tests, and fault locations. We mine bug reports from issue tracker links provided by Defects4J. For 18 cases from the Chart project without a clear bug issue link, we use the fixing commit message as the bug description, and to prevent potential data leakage, we manually verified these commit messages to ensure they do not contain explicit fix instructions. Similar to the study (Nashid et al., 2025), we collect the failing tests provided by Defects4J. Consistent with common practices in the field of program repair (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we assume perfect fault localization and provide fault localization context, which is derived from the developer’s bug fix (following the process in Section 3.2), directly into the agent’s context. Specifically, similar to prior work (Bouzenia et al., 2025), this context consists of all file(s) and line(s) that need to be edited to fix the bug.
Historical data collection.
Motivated by the preliminary study in Section 3, which shows that 71.1% of bugs are blameable and most map to a single unique blame commit, we collect history per bug as follows:
  • Blameable with one unique commit: Run git blame on deleted or modified lines to obtain the blame commit, as defined in Section 4.1.3.
  • Blameable with multiple commits: Only three bugs fall in this case, we use the LLM-as-a-judge (the same LLM as HAFixAgent) to select the single most relevant commit. The model is prompted with the set of blameable lines and asked to choose the line which is most likely related to the bug’s root cause.
  • Blameless with zero commits: Apply the nearest line fallback strategy from Section 4.1.3 to select a blame commit.
Given the selected blame commit, we extract and construct data of fn_all, fn_pair, and fl_diff as defined in Section 4.1.3.
5.2. Baselines
We compare HAFixAgent with two recent APR techniques as baselines, i.e., RepairAgent (Bouzenia et al., 2025) and BIRCH-feedback (Nashid et al., 2025).
RepairAgent.
RepairAgent is an autonomous LLM-based agent equipped with pre-defined API tooling. The original study evaluated it on 835 bugs (from Defects4J v1.2 and v2), reporting 186 plausible (test-passing) and 164 correct (manual semantic check) fixes. Our comparisons are performed on the 829-bug overlap with our v3 set. For our aggregate effectiveness comparison, we compare our test-passing patches against their reported 186 plausible fixes, as these metrics are equivalent. For our fine-grained, per-category analysis, we are limited to their 164-bug correct set, as bug IDs (which are required for categorization) were only released for this set instead of the plausible bug set. [^2] https://github.com/sola-st/RepairAgent/issues/18
BIRCH-feedback.
BIRCH-feedback targets multi-hunk repairs. We compare results on the 371 overlapping multi-hunk bugs reported in that study and use their best-performing configuration with feedback as the baseline, which we refer to as BIRCH-feedback.
5.3. Metrics
In line with previous studies (Chen et al., 2021b; Parasaram et al., 2025; Ehsani et al., 2025; Nashid et al., 2025; Silva and Monperrus, 2025), we report Plausible@1 to evaluate the correctness based on whether the generated patch passes all test cases. Formally:
(1) $\mathrm{Plausible@1}=\frac{1}{|\mathcal{B}|}\sum_{b\in\mathcal{B}}\mathrm{TestPass}\!\left(\text{patch}^{\,b}_{1}\right)$
where $\mathcal{B}$ denotes the set of bugs and $\mathrm{TestPass}(\cdot)\in\{0,1\}$ returns $1$ if the patch compiles and all tests pass, and 0 otherwise. We generate one patch per bug using standard decoding settings, with the temperature set to 0.0 for enhancing reproducibility.
Complementing Plausible@1, which measures the pass rate as a proportion of all bugs), we also report #Pass (the absolute count of bugs passing the test suite), similar to prior work (Nashid et al., 2025). To quantify the unique contribution of diverse historical context, we additionally report #Unique Pass, to show the number of bugs solved by a configuration using historical context but not by its non-history counterpart. We visualize overlaps using Venn diagrams to illustrate multi-way complementarity.
For efficiency, we report the average number of agent steps per bug, the average cost per bug in USD computed from token usage, and the cost stratified by outcome (successful versus failed fixes). To assess the statistical significance of differences in cost and number of steps, we compare the distributions across the four context configurations. Since the data is paired (same bugs, different configurations) and not assumed to be normally distributed, we employ the Friedman test (Friedman, 1937), a non-parametric test for multiple comparisons. If the Friedman test is significant ($p<0.05$), we conduct pairwise Wilcoxon signed-rank tests (Wilcoxon, 1992) for post-hoc analysis to compare each history-aware configuration against the non-history configuration. To control for multiple comparisons (three pairs), we apply the Bonferroni correction (Dunn, 1961) and set the significance threshold $\alpha$.
5.4. Implementation
HAFixAgent is implemented in Python and builds on mini-swe-agent (Yang et al., 2024) for the execution loop, which aligns with our lightweight design. While mini-swe-agent (and swe-agent) is designed naturally to support solving PR issue benchmarks such as SWE-Bench (Jimenez et al., 2024), it lacks the capability of historical context management (e.g., history localizing and retrieval), the repairing strategies and tools specifically for resolving Defects4J bugs. It is an open-source, lightweight framework with about 100 lines of core loop code, enabling seamless integration of historical context extraction and bash tools orchestration during runtime. We use DeepSeek-V3.2-Exp (DeepSeek-AI et al., 2024; DeepSeek-AI, 2025) via the official API with temperature set to 0.0, based on the official DeepSeek recommendation (DeepSeek, 2025). Prompts are rendered at runtime with jinja2.
We run HAFixAgent on Ubuntu 20.04, with each bug executing in an isolated Docker container. This sandbox environment is built from the released Defects4J image. For this evaluation, we implement the abstract compile and test tools (described in Section 4.3) as wrappers for the defects4j compile and defects4j test -r commands, respectively. The container is preinstalled with all necessary Defects4J and standard bash tools required by HAFixAgent. For each bug, we cap the agent’s loop at 50 steps, enforce a per-bug cost guard of \$1 USD, and set a 1-hour timeout. After each run, the container is automatically cleaned, while all generated patches and logs are persisted for analysis.
6. Results
In this section, we provide the motivation, approach, and results for each of our research questions.
6.1. RQ1: How Effective Is HAFixAgent at Fixing Real-World Bugs?
6.1.1. Motivation
Our prior work (Shi et al., 2025) shows that injecting historical context improves LLM repair for single-line bugs under simple prompting. History carries heuristic signals that are hard to infer from a static snapshot: the fault introducing change, coevolving functions and files, and often intent hints in commit diffs and messages. RQ0’s finding that most Defects4J bugs, even multi-hunk ones, have only one single historical blame commit, suggests that historical context could also scale to APR techniques for more complex bug types.
This RQ rigorously validates the benefit of blame-derived context on SH, SFMH and MFMH bugs, in the context of our proposed HAFixAgent. To thoroughly validate our approach, we first establish HAFixAgent’s practical viability by benchmarking it against state-of-the-art (SOTA) baselines to determine if our history-aware method is competitive. Second, we isolate the specific impact of our core hypothesis by evaluating the agent across different context configurations (i.e., ”with” versus ”without” history). This analysis is essential to pinpoint and quantify the precise contribution of historical context and to understand how these gains vary by bug category. This motivates the following questions: (a) is HAFixAgent competitive against SOTA baselines? (b) what is the performance impact of adding historical context to an APR agent, and (c) how do these gains vary by bug category?
6.1.2. Approach
We compare HAFixAgent against two SOTA baselines on shared bug subsets: RepairAgent (829 bugs) and BIRCH-feedback (371 multi-hunk bugs). For RepairAgent, our direct comparison uses the set of bugs they reported as correctly patched, as their full list of plausible bug IDs is unavailable; we also report their published aggregated counts (Section 5.2) for reference. To analyze performance across different bug types, we categorize results by our four bug types (SL, SH, SFMH, MFMH). We pay special attention to the multi-hunk categories (SFMH and MFMH), as their need for coordinated changes may make them more sensitive to historical context. BIRCH-feedback serves as a strong baseline for this specific subset.
In addition, we evaluate the four context configurations defined in Section 4.1 across all 854 Defects4J bugs: HAFixAgent-non-history, HAFixAgent-fn_all, HAFixAgent-fn_pair, and HAFixAgent-fl_diff. Notably, 247 bugs (28.9%) are blameless (fix adds only new lines), precluding standard blame. To enable these blameless bugs to also benefit from historical context, we apply the fallback strategy (defined in Section 4.1.3), ensuring every bug has meaningful historical context.
All experiments use the same LLM, identical parameters, and a consistent containerized runtime. We report both Plausible@1 (computed via termination checks in Section 4.2) and #Pass (the total number of bugs with patches passing all tests). To quantify the specific benefit of historical context, we also report #Unique Pass for the number of bugs repaired by a history-aware configuration that were not repaired by the non-history configuration.
6.1.3. Results
Table 3. Effectiveness comparison of HAFixAgent against RepairAgent and BIRCH-feedback on their shared subsets. For HAFixAgent and BIRCH-feedback, cells report the number of bugs with plausible patches that pass the full test suite. For RepairAgent, we report the number of bugs with correct patches because their repository does not provide plausible bug IDs, and only reports aggregates (plausible = 186, correct = 164). Since correct is a subset of plausible, RepairAgent’s counts are more conservative compared to the plausible numbers for BIRCH-feedback.
(a) 829 common bugs with RepairAgent
Category Common HAFixAgent RepairAgent
non-history fn_all fn_pair fl_diff
SL 165 135 134 134 138 99
SH 128 93 91 102 98 30
SFMH 408 224 238 235 239 31
MFMH 128 50 47 43 48 4
Total 829 502 510 514 523 164 (186*)
    The 186* denotes the total number of plausible patches of RepairAgent; specific bug IDs are only available for correct patches in their repository.
(b) 371 common multi-hunk bugs with BIRCH-feedback
Category Common HAFixAgent BIRCH-feedback
non-history fn_all fn_pair fl_diff
SFMH 244 123 125 133 126 100
MFMH 127 49 46 42 47 33
Total 371 172 171 175 173 133
HAFixAgent substantially outperforms both RepairAgent and BIRCH-feedback. On the 829 bugs shared with RepairAgent (Table 3(a)), the best context configuration of HAFixAgent (fl_diff) mode repairs 523 bugs versus 164 correct fixes reported by RepairAgent, a +218.9% overall improvement. Across different context configurations of HAFixAgent, the improvements relative to this 164-fix baseline range from +206.1% to +218.9% (502, 510, 514, 523), with an average of 212.3% improvement. When comparing against RepairAgent’s reported aggregate plausible total of 186 (the equivalent metric), the corresponding improvements are +170.0% to +181.2% (175.4% by average).
On the 371 bugs shared with BIRCH-feedback (Table 3(b)), HAFixAgent also achieves higher totals. The best context configuration (fn_pair) repairs 175 bugs versus 133 for BIRCH-feedback, a +31.6% improvement. Across four context configurations, the gains range from +28.6% to +31.6%, with an average of 29.9% improvement. By category, the best HAFixAgent configuration (fn_pair) repairs 133 SFMH and 42 MFMH bugs, compared with 100 and 33 for BIRCH-feedback. Overall, HAFixAgent is consistently effective than both SOTA baselines across shared subsets and categories.
We note, however, that these SOTA comparisons are confounded by the use of different LLMs (RepairAgent used GPT-3.5, BIRCH-feedback used o4-mini, and our work uses DeepSeek-V3.2-Exp). Therefore, this analysis primarily establishes HAFixAgent’s competitiveness, while our internal ablation study (comparing history to non-history with the same LLM) provides the direct, controlled evidence for our claims.
Table 4. Repair performance across different context configurations over 4 bug categories on Defects4J. # Pass represents the number of bugs passing the test suite. Unique # Pass represents the number of bugs solved by that configuration but not by the corresponding non-history configuration. The highest value for each category group in each column is highlighted in bold.
Category Context # Pass Plausible@1 # Unique Pass
SL non-history 137 82.0% -
fn_all 136 81.4% 10
fn_pair 136 81.4% 10
fl_diff 140 83.8% 11
SH non-history 97 72.4% -
fn_all 96 71.6% 16
fn_pair 108 80.6% 19
fl_diff 103 76.9% 18
SFMH non-history 238 56.0% -
fn_all 251 59.1% 67
fn_pair 249 58.6% 69
fl_diff 254 59.8% 68
MFMH non-history 50 39.1% -
fn_all 47 36.7% 13
fn_pair 43 33.6% 13
fl_diff 48 37.5% 12
Figure 3. Venn diagram comparing bug repair performance across 4 context configurations (non-history, fn_all, fn_pair, and fl_diff)
over 4 bug categories on Defects4J.
History configurations fix more unique bugs compared to the non-history configuration, with the largest gains for the multi-hunk bug categories. Table 4 summarizes the number of passing and uniquely passing bugs across the four context modes, while Figure 3 visualizes their overlap and complementarity. Across all 854 bugs, the three history-based configurations add 194 bugs that the non-history configuration never fixes, whereas the non-history configuration contributes only 32 unique fixes not recovered by any history configuration. By category, fl_diff performs best on SL with 140 total fixes and 11 unique ones and fn_pair leads on SH with 108 fixes and 80.6% Plausible@1. In SFMH, fl_diff achieves 254 fixes and the three history configurations collectively add 119 more repairs, while the non-history configuration uniquely fixes 17. Finally, the MFMH category reveals a complex trade-off. Although the non-history configuration achieves the highest total fixes (50), this aggregate score masks a clear complementarity. Collectively, the history configurations add 24 unique fixes not found by the non-history agent, outweighing the 8 unique fixes from the non-history side (Figure 3). This suggests that while history can sometimes distract the agent on complex multi-file bugs, it also provides an orthogonal signal that unlocks repairs for a different subset of bugs.
The three history configurations are complementary rather than redundant, solving distinct sets of bugs that others miss. The Venn diagrams in Figure 3 show large shared cores in each category (SL 118, SH 72, SFMH 140, MFMH 22), yet each heuristic contributes nontrivial unique fixes, especially in SFMH, where fn_all, fn_pair, and fl_diff add 21, 19, and 19 unique cases, respectively. This pattern shows that fn_all captures co-evolution across functions, fn_pair captures before and after semantics, and fl_diff captures fine-grained edits at the file level, so the union of history-based configurations yields broader repair coverage than any single configuration.
Summary for RQ1: (1) HAFixAgent outperforms RepairAgent by an average of 212.3% across the different context configurations, and outperforms BIRCH-feedback by an average of 29.9%. Additionally, across all 854 bugs, history configurations add 194 history-only fixes versus 32 non-history-only, with the largest gains in multi-hunk categories. (2) The three history heuristics solve distinct cases beyond a shared core, indicating that combining heuristics yields broader coverage than any single setting.
6.2. RQ2: What Are the Cost and Efficiency of HAFixAgent?
6.2.1. Motivation
While RQ1 demonstrates that incorporating historical information into HAFixAgent enhances repair performance, it is also important to understand the associated cost and efficiency trade-offs. Cost is a first-order concern for agent systems because iterative tool use can accumulate substantial token and time budgets, even on simple bugs (Bouzenia et al., 2025; Xia and Zhang, 2024). Recent evaluations also highlight the risk of “token snowballs” and expensive failures (a phenomenon where unresolved attempts consume several times more tokens than successful ones (Fan et al., 2025), making cost and efficiency as important to measure as accuracy. HAFixAgent integrates additional contextual information into each prompt, which might potentially increase inference cost and agent reasoning steps. Therefore, we analyze whether the performance gains observed in RQ2 come at a significant cost and how efficiently the agent converges toward successful repairs across bug categories.
6.2.2. Approach
We evaluate the inference cost (USD) and agent reasoning steps across all four configurations (non-history, fn_all, fn_pair, and fl_diff) over four bug categories (SL, SH, SFMH and MFMH). Inference cost is computed from input and output token usage converted to USD using official API pricing rates published for the DeepSeek-V3.2-Exp model. Efficiency is measured by the number of agent actions taken per run. A run terminates either upon a successful repair or by hitting one of the predefined limits (50 steps, \$1 USD cost, or 1-hour timeout), as defined in Section 5.4. We report cost and step metrics separately for successful and failed attempts, allowing us to isolate and analyze the average resources required for a successful repair. Finally, we analyze the success rate versus cost across configurations to characterize category-specific trade-off frontiers.
6.2.3. Results
(a) SL
(b) SH
(c) SFMH
(d) MFMH
Median values (Success / Failed)
SL SH SFMH MFMH
non-history 12 / 50 23 / 50 24 / 50 32 / 50
fn_all 14 / 50 25 / 50 27 / 50 28 / 50
fn_pair 13 / 50 24 / 50 25 / 50 31 / 50
fl_diff 16 / 50 22 / 50 29 / 50 29 / 50
Figure 4. Agent reasoning step distributions for successful and failed repairs across different HAFixAgent context configurations.
(a) SL
(b) SH
(c) SFMH
(d) MFMH
Median values (Success / Failed)
SL SH SFMH MFMH
non-history 0.005 / 0.038 0.016 / 0.047 0.016 / 0.045 0.029 / 0.038
fn_all 0.008 / 0.050 0.019 / 0.046 0.020 / 0.047 0.026 / 0.047
fn_pair 0.005 / 0.046 0.015 / 0.039 0.016 / 0.045 0.023 / 0.040
fl_diff 0.008 / 0.042 0.014 / 0.047 0.021 / 0.047 0.023 / 0.043
Figure 5. Inference cost distributions for successful and failed repairs across different HAFixAgent context configurations.
Successful repairs converge well before the step cap. Figure 4 reports agent step distributions by configuration and outcome. Across categories, the median number of steps for successful repairs falls far below 50, whereas for failed attempts it sits at the cap of 50 for all configurations. The median number of steps for successful repairs is 12-16 in SL, 22-25 in SH, 24-29 in SFMH, and 28-32 in MFMH, showing that harder categories require more steps but still finish well under the cap.
History does not significantly increase step counts. To rigorously assess whether historical heuristics significantly alter repair costs or the number of steps compared to the non-history configuration, we conducted Friedman tests followed by pairwise Wilcoxon signed-rank tests with Bonferroni correction ($\alpha=0.0167$) on bugs successfully fixed by all four configurations (Table 5). Within each category, the success medians for non-history, fn_all, fn_pair, and fl_diff differ by only a few steps. The Friedman tests on this matched set detect no significant step differences in any category (all $p\geq 0.05$). Notably, in MFMH, the history configurations are as fast or faster than the non-history configuration (non-history 32 vs fn_all 28, fn_pair 31, fl_diff 29), and in SH, the fl_diff configuration converges the quickest (median 22). This confirms that adding historical context does not create a significant step overhead.
Agent cost is dominated by failure, not configuration. Figure 5 summarizes inference cost distributions by configuration and outcome. Across categories, failed attempts cost markedly more than successful ones, clustering near 0.03-0.05 USD median, while successes are cheaper: SL 0.005-0.008, SH 0.014-0.019, SFMH 0.016-0.021, and MFMH 0.023-0.029. This mirrors the step analysis, where failures hit the step cap and thus consume more tokens.
Adding historical context keeps the cost comparable in most configurations. For SL, the Friedman test is significant for cost, with the Wilcoxon post-hoc tests showing that fn_all ($p=0.0022$) and fl_diff ($p=0.001$) cost more than non-history, while fn_pair shows no significant difference ($p=0.0444>0.0167$). For SH, SFMH, and MFMH, Friedman tests are not significant for cost (all $p\geq 0.05$). Notably, the median costs for successful repairs (shown in Figure 5) can be as low or lower with history in harder categories (for example, MFMH median success costs drop from 0.029 to 0.026-0.023), but these reductions are not statistically significant on the matched set (i.e., the subset of bugs fixed by all four configurations).
Figure 6. Trade-off analysis of success rate and cost per bug across different combinations of configurations.
Table 5. Statistical Comparison: History Heuristics vs. Non-history (Bonferroni $\alpha=0.0167$). The last three columns report p-values for each history configuration compared against the non-history configuration.
Category Metric N Friedman p Pairwise p-values
fn_all fn_pair fl_diff
SL Cost 118 <0.001<0.001 0.0022 0.0444 0.001
Steps 118 0.176
SH Cost 72 0.184
Steps 72 0.400
SFMH Cost 140 0.122
Steps 140 0.051
MFMH Cost 22 0.142
Steps 22 0.427
  • N = number of bugs successfully fixed by all 4 configurations.
  • Bold values indicate $p<0.0167$.
  • — indicates the Friedman test was not significant ($p\geq 0.05$); post-hoc tests were not performed.
Combining history configurations increases success but also increases cost, exhibiting diminishing returns, and the most cost-effective choices are usually two or three configurations. Figure 6 plots success rate versus average cost per bug by category, with marker shape indicating how many configurations are combined. For Single-Line bugs, fn_pair +fl_diff reaches 90.4% at an average cost of 0.041 USD, close to the union of all four configurations, 95.2% at 0.086 USD. For Single-Hunk bugs, fn_all +fn_pair attains 87.3% at 0.063, while the four-configuration union gives 94.0% at 0.130. For Single-File-Multi-Hunk bugs, fn_all +fl_diff yields 73.9% at 0.074, compared to 84.0% at 0.147 with the four-configuration union. For Multi-File-Multi-Hunk bugs, non-history +fn_pair provides 49.2% at 0.086 before the cost rises to 57.8% at 0.182 with all four. Across categories, two or three history configurations give the best success per dollar. As these results show, adding the final configuration to create the four-configuration union raises the cost substantially for only small gains, demonstrating clear diminishing returns.
Summary for RQ2: (1) Successful repairs finish well before the step cap, while failures commonly hit the cap. (2) History-based APR agents are as efficient as non-history in terms of step count and cost (except for SL). (3) Combining two or three history configurations delivers the best success, while using all configurations yields small extra gains at a significantly higher cost.
7. Discussion and Future Work
In this section, we provide a discussion of our results and introduce future work.
7.1. Performance Regressions
We observe 32 cases where a history configuration underperforms the non-history one or misses non-history fixes (Figure 3). For instance, in MFMH, the non-history configuration attains 50 while fn_all, fn_pair, and fl_diff reach 47, 43, and 48, respectively, even though each contributes 12 to 13 unique fixes that the non-history does not solve. This pattern aligns with LLM evidence that more context can hurt: models are easily distracted by irrelevant additions (Shi et al., 2023), and performance often declines when relevant facts sit in the middle or deeper parts of long prompts (Levy et al., 2024). In our setting, extra historical snippets can dilute salient context and reduce Plausible@1 on some bugs that don’t need historical context. Future work might filter and summarize history for direct relevance or add a guardrail that falls back to a leaner prompt when early steps show no progress.
7.2. Heuristics Complementarity
A key finding from RQ1 (Figure 3) is that the three historical heuristics are complementary, not redundant. Each historical heuristic (e.g., fn_all, fn_pair, fl_diff) repairs a distinct set of bugs that the others miss. This has important implications for why history is effective. It suggests that different bug types are susceptible to different forms of historical context:
  • fl_diff (file-level diff), the strongest overall, likely excels at capturing fine-grained and localized edit patterns that are common in single-line and single-hunk bugs.
  • fn_pair (historical function before and after) provides a richer semantic context, showing the agent how a function’s logic was intended to operate, which may be crucial for more complex logic errors.
  • fn_all (all functions from blame commit) captures co-evolutionary patterns (e.g., when function A changed, function B also changed), which is a vital clue for multi-hunk bugs where a fix requires coordinated edits.
This complementarity also explains the practical trade-off identified in RQ2 (Figure 6): combining two or three heuristics shows a higher pass rate and yields the most cost-effective performance, as it provides the agent with a diverse set of contexts.
7.3. Future Work
Richer historical heuristics beyond the blamed commit.
HAFixAgent currently injects three compact heuristics drawn from the single commit that last touched the buggy lines. A broader temporal view may help: commits immediately before or after the blamed change, refactorings, and coevolution bursts can reveal the mechanism that produced the bug. Future work can build coevolution-aware retrieval over historical commit windows, summarize change intents such as API migrations or dependency upgrades, and learn when to expand beyond one commit. Such a broader historical context may improve robustness, particularly for bugs whose causes span multiple commits.
Enhanced context for blameless bugs.
Building on our discussion in Section 4.1.3, the simple nearest-line fallback for the 28.9% of blameless bugs can be significantly improved. Instead of just finding the nearest line, a future agent could analyze the history of the parent class or module to infer project-specific patterns for adding new functionality. It could also search the repository for semantically similar add-only commits (e.g., how has this project added null-checks in other files?) to provide the agent with a template for the kind of code it needs to write.
Assessing history’s impact on patch quality.
A valuable future direction is to move beyond test-pass-based evaluation to investigate how historical context impacts the quality and maintainability of the generated patches. While multiple agent configurations might produce a plausible fix, it is an open question whether the history-aware patches are qualitatively superior. For instance, a patch generated with historical context might be more idiomatic, such as re-using a specific helper function or logging style from the historical diff, and more semantically aligned with the project’s evolution. This could contrast with a non-history patch that, while functional, acts as a brute-force fix that simply passes the test suite. Future work could investigate this by designing automated tools to rate patch quality or by employing static analysis metrics to measure code complexity and semantic differences.
Extending to a multi-agent repair system.
A promising, though more complex, frontier is to evolve HAFixAgent from a single-agent system into a collaborative multi-agent architecture (Rafi et al., 2024; Lin et al., 2025; Lee et al., 2024). Instead of one agent handling all tasks, we could design a system with specialized roles. For instance, a Coordinator agent could oversee the process, dispatching a specialized History agent to analyze repository patterns, blame commits, and issue reports to form a repair hypothesis. This hypothesis would then be passed to a Fixer agent focused on patch generation. Critically, a Reviewer agent could then critique the proposed patch, checking it not only against the test suite but also against the historical context for style, conventions, and potential side effects, which is similar to human code review. This separation of concerns could be particularly effective for the complex multi-file-multi-hunk (MFMH) bugs, where a single agent’s reasoning chain and context window can be overwhelmed.
Applying historical context to other SE tasks beyond APR
The same history-based heuristics may benefit other SE code tasks such as code generation, code completion, code review, refactoring, and test generation. Repository history heuristics such as coevolution patterns, change intent, and dependency updates can guide model choices and keep outputs aligned with project conventions. Future work should evaluate these heuristics on tasks beyond repair using task-specific metrics (functional correctness, style consistency, temporal consistency, and safety), compare against retrieval baselines, and study generalization across languages and repositories.
8. Threats to Validity
In this section, we discuss the threats to validity of our study about HAFixAgent.
8.1. Internal Validity
Data leakage.
DeepSeek-V3.2-Exp is available as open weights, but the pretraining corpus for this model is not disclosed to the best of our knowledge. As a result, we cannot verify if the specific projects or bugs in Defects4J are included in the pretraining data. However, our evaluation of HAFixAgent focuses on the relative effectiveness improvement of different historical heuristics, rather than their absolute performance. For instance, if HAFixAgent-fl_diff fixes a bug that the HAFixAgent-non-history does not fix, this improvement is attributed to the benefit of historical context, regardless of whether the LLM has seen the bug before. Our evaluation design naturally isolates and mitigates the possible influence of potential data leakage. To fully resolve this threat, we would need to retrain DeepSeek-V3.2-Exp from scratch, which would be infeasible for an academic project.
Fault localization.
Our evaluation assumes perfect fault localization to isolate the evaluation of historical context on patch generation, a common setup in LLM-based APR studies (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025). Note that even given perfect fault localization, as most of the current work did, the LLM-based APR still struggle with complicated bugs (e.g., 39.1% Plausible for non-history configuration in MFMH bugs in our study). In practice, FL is noisy, especially for human-reported issues and multi-hunk defects, which can lower repair rates and lead to potentially wrong repair directions. For iterative agentic APR, updating FL during repair might help, but it does not eliminate the challenge. So we leverage perfect fault localization to isolate the evaluation of the historical context in HAFixAgent.
8.2. External Validity
Dataset and language generalizability.
Our study uses Defects4J v3.0.1 with 854 Java bugs. Results may not transfer to other ecosystems where language, project size, test adequacy, and repository practices differ. The history profile we observe in Defects4J (for example, many bugs mapping to a single history commit) can change in other benchmarks, repositories with frequent code moves, or workflows that rewrite history. To mitigate this threat, we expose a dataset-agnostic interface behind the history extractor and provide dataset adapters that can be subclassed when porting to languages such as Python or C. We release code and scripts so other communities can rerun the pipeline on local datasets.
LLM and agent generalizability.
HAFixAgent is instantiated with one agent loop, a narrow bash tool set, specific prompts, and a single LLM configuration with fixed step, cost, and time guards. Different agents, tool scopes, models, context windows, or decoding policies may change absolute numbers and the relative strength of the three history heuristics. API level determinism does not guarantee identical runs in practice, so small variation across executions and providers is expected. To reduce this threat, we use DeepSeek-V3.2-Exp as a recently released strong model at the time of writing. We also assume perfect fault localization to isolate the effect of history, and generalizing to noisy localization remains future work.
9. Related Work
In this section, we discuss related work about the traditional automated program repair, LLM and agent based automated program repair, and in-context learning for automated program repair.
9.1. Traditional Automated Program Repair
Traditional APR formulates patch generation as a search or constraint problem over program transformations guided by tests or specifications. Search-based systems evolve patches with genetic or heuristic search over mutation operators and validate candidates against the test suite, exemplified by GenProg and successors (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018). Semantics or constraint-based repair derives patches by solving synthesis constraints obtained from symbolic execution or program analysis, as in SemFix (Nguyen et al., 2013) and Angelix (Mechtaev et al., 2016), and with targeted condition synthesis such as Nopol (Xuan et al., 2017). Template-based repair uses human-designed or mined fix schema that match buggy contexts and instantiate edits such as TBar (Liu et al., 2019b) and AVATAR (Liu et al., 2019a). A complementary line reuses in-project code fragments as ingredients, such as ssFix (Xin and Reiss, 2017) and CAPGEN (Wen et al., 2018). These approaches established core pipelines for localization, candidate generation, and test-based validation, but their coverage is constrained by operator or pattern design and by search scalability.
Learning-based APR before LLMs frames repair as code-to-code translation or edit prediction trained on a large amount of code corpus. Early neural systems include DeepFix (Gupta et al., 2017) for syntax errors and sequence-to-sequence repair, such as SequenceR (Chen et al., 2021a), with later improvements from model ensembling and syntax guidance in CoCoNuT (Lutellier et al., 2020), CURE (Jiang et al., 2021), and Recoder (Zhu et al., 2021). Template mining and ranking were also brought to production, for example, Facebook’s Getafix (Bader et al., 2019) for static-analyzer warnings and the end-to-end SapFix (Marginean et al., 2019) pipeline that proposed developer-reviewed patches. Together, traditional APR demonstrates effective repair across bug classes while revealing recurring limitations in computation cost, low accuracy and overfitting to tests. These limitations have motivated LLM and agent-based APR that learn fixes from data and plan tool-driven repair loops, which we introduce next.
9.2. LLM and Agent based Automated Program Repair
Large language models (LLMs) have significantly advanced Automated Program Repair (APR), moving from direct prompting to sophisticated, multi-step agentic systems. Early evidence shows code LLMs already outperform prior learning-based APR and benefit from fine-tuning and careful evaluation on standard benchmarks (Jiang et al., 2023; Xia et al., 2023b). This led to a variety of prompting and interaction strategies. For example, conversation-driven repair simulates a developer’s debugging process by interleaving patch generation with test feedback, achieving a strong balance of cost and accuracy (Xia and Zhang, 2024). Other approaches focus on improving the model’s reasoning, such as self-directed repair, which uses a chain-of-thought process to gather knowledge before fixing (Yin et al., 2024). The plastic surgery line revisits ingredient reuse by aligning LLMs with project-specific code (Xia et al., 2023a). Beyond pure prompting, hybrid and template-guided methods constrain or scaffold the patch space with analysis or repair templates to improve plausibility and correctness (Li et al., 2025a; Huang et al., 2025).
Agent-based APR systems represent a further step, shifting from single-shot patch generation to autonomous, tool-using workflows. Pioneering general-purpose software agents, like SWE-agent (Yang et al., 2024) and OpenHands (Wang et al., 2025), established a powerful paradigm by equipping LLMs with tools to interact with a repository, such as editors, shell commands, and test runners. Building on this, repair-focused agents have integrated more specialized tools. For instance, some employ search and fault localization to narrow down the buggy code (Bouzenia et al., 2025; Zhang et al., 2024c), while others explore multi-agent collaboration to divide the debugging task (Lee et al., 2024), with empirical studies beginning to map this rapidly evolving design space (Meng et al., 2024). A recent trend also focuses on improving agent performance on complex, repository-level tasks by incorporating memory and experience, allowing agents to learn from prior repair trajectories (Mu et al., 2025; Chen et al., 2025). These lines generally rely on local code and runtime outputs, with limited study of commit history as a first-class context inside the loop, which is the focus of our work.
9.3. In-context Learning for Automated Program Repair
A core question for LLM-based APR is which context to provide and how to structure it. Prior work shows that adding bug-related facts such as error messages, stack traces, and failing tests improves repair, and that the choice of facts matters (Parasaram et al., 2025; Xia and Zhang, 2024). Studies on local context indicate that models are sensitive to how much and which code surrounds the edit region (Prenner and Robbes, 2024). Other work leverages natural language bug reports or transforms descriptions into edits, and explores execution traces to guide patching (Fakhoury et al., 2024; Haque et al., 2025). Repository-level evaluations report that larger, realistic tasks remain challenging and require broader context curation (Chen et al., 2024).
Beyond local signals, historical context is receiving attention. HAFix (Shi et al., 2025) demonstrates that blame-derived commit data is an effective, lightweight historical signal for fixing more bugs, motivating history-aware agent designs at Defects4J scale. Layered knowledge injection systematically adds project and repository knowledge to prompts and improves bug fixing on broader-context scenarios (Ehsani et al., 2025). Built on top of HAFix, HAFixAgent advances this direction by bringing blame-driven history into the agentic workflow and quantifying its effect on single-hunk and multi-hunk bugs.
10. Conclusion
This paper presented HAFixAgent, a history-aware agentic approach that injects blame-derived repository context into an observe-act-verify loop to address the more complex single- and multi-hunk bug categories. A preliminary study shows how blame commit history is not only widely available (71.1% blameable), but also is surprisingly concentrated with 70.7% of Defects4J bugs having exactly one unique blame commit across their buggy line.
When applying HaFixAgent on all 854 Defects4J bugs, history-aware settings were able to add 194 history-only fixes that the non-history configuration never achieves. When compared to SOTA baselines, HAFixAgent outperforms RepairAgent by an average of 212.3% and BIRCH feedback by 29.9%. Among the three historical contexts, fl_diff is the strongest overall and fn_pair leads on single-hunk bugs. Successful runs finish well before the 50-step cap, and the median token cost does not increase; in the hardest multi-file-multi-hunk category, costs are lower than the non-history configuration. Lastly, combining two or three heuristics captures most gains, while leveraging all heuristics adds little benefit for a significantly higher cost. These findings give a simple recipe for practice: (1) ground the agent loop in version control history, (2) prefer diff-based context, (3) add one complementary heuristic when needed, and (4) avoid collecting excessive context.
This study has natural limitations. We assume perfect fault localization (similar to related work), focus on one benchmark and language, and use a single strong agent model and loop. Future work will relax these assumptions by evaluating additional datasets and languages, testing imperfect localization, learning to select or combine history heuristics at runtime, and integrating richer project artifacts. We expect progressive use of historical context to remain a robust and lightweight direction for advancing agent-based program repair.
Appendix A HAFixAgent Prompt
System Prompt of HAFixAgent. The values inside placeholder will be rendered in runtime.