### v3

# HAFixAgent: History-Aware Program Repair Agent

Abstract.
Automated program repair (APR) has recently shifted toward large language models and agent-based systems, yet most systems rely on local snapshot context, overlooking repository history. Prior work shows that repository history helps repair single-line bugs, since the last commit touching the buggy line is often the bug-introducing one. In this paper, we investigate whether repository history can also improve agentic APR systems at scale, especially for complex multi-hunk bugs. We present HAFixAgent, a History-Aware Bug-Fixing Agent that injects blame-derived repository heuristics into its repair loop. A preliminary study on 854 Defects4J (Java) and 501 BugsInPy (Python) bugs motivates our design, showing that bug-relevant history is widely available across both benchmarks. Using the same LLM (DeepSeek-V3.2-Exp) for all experiments, including replicated baselines, we show: (1) Effectiveness: HAFixAgent outperforms RepairAgent (+56.6%) and BIRCH-feedback (+47.1%) on Defects4J. Historical context further improves repair by +4.4% on Defects4J and +38.6% on BugsInPy, especially on single-file multi-hunk (SFMH) bugs. (2) Robustness: under noisy fault localization (+1/+3/+5 line shifts), history provides increasing resilience, maintaining 40 to 56% success on SFMH bugs where the non-history baseline collapses to 0%. (3) Efficiency: history does not significantly increase agent steps or token costs on either benchmark.
Automated Program Repair, Agentic Software Engineering, Large Language Model, Mining Software Repository, Software Engineering Agent, AI Agent, Bug Fixing
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software testing and debuggingccs: Computing methodologies Artificial intelligence
1. Introduction
Software bugs are an inevitable and costly aspect of software development, leading to system failures, security vulnerabilities, and degraded user experience (BOULDER, 2013; Zhang et al., 2024a; Weiß et al., 2007). Since manually identifying and fixing these bugs consumes significant developer time and resources, automated program repair (APR) has been an active and promising research direction for over a decade, aiming to mitigate this burden by automatically generating patches for detected bugs (Le Goues et al., 2019). APR has evolved from heuristic-based (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018), constraint-based (Le et al., 2017; Mechtaev et al., 2016; Long and Rinard, 2015), and template-based techniques (Liu et al., 2019b; Martinez and Monperrus, 2016; Liu et al., 2019a) to learning-based methods (Chen et al., 2021b; Jiang et al., 2021; Li et al., 2020; Lutellier et al., 2020; Zhu et al., 2021, 2023; Ye and Monperrus, 2024) and, most recently, large language model-based approaches (Zhang et al., 2024b; Jiang et al., 2023; Hossain et al., 2024; Xia et al., 2023b; Xia and Zhang, 2024; Xia et al., 2023a; Fan et al., 2023; Yin et al., 2024; Huang et al., 2025; Li et al., 2025a; Shi et al., 2025) and agent-based approaches (Yang et al., 2024; Bouzenia et al., 2025; Mu et al., 2025; Meng et al., 2024; Lee et al., 2024; Zhang et al., 2024c; Chen et al., 2025; Wang et al., 2025). However, despite substantial progress, reliably repairing complex bugs remains a persistent challenge.
Particularly difficult are bugs that span multiple lines and locations (e.g., multi-hunks), which require coordinated, non-local edits. As Nashid et al. (2025) highlight, these bugs are common yet significantly harder for current LLM-based approaches.
To address such complex bugs, recent APR advances have pursued two promising yet separate directions. First, agent-based approaches (Bouzenia et al., 2025; Yang et al., 2024; Wang et al., 2025) leverage LLMs with iterative tool use (code search, editing, testing) to tackle complex reasoning, but typically operate only on the current code snapshot without deeper historical context. Second, drawing from decades of mining software repositories (MSR) research (Sliwerski et al., 2005; Hassan, 2006; Adams et al., 2010), prior work (Shi et al., 2025) showed that injecting historical context derived from git blame [^1] https://git-scm.com/docs/git-blame commits can improve LLM repair for single-line bugs, but lacked the iterative reasoning capabilities of agents.
We hypothesize that combining agentic workflows with repository history can significantly enhance APR, particularly for complex multi-hunk bugs. History provides information on relevant past changes, developer rationale, and co-evolution patterns, while the agent leverages tools to explore and validate repairs. Yet, it is unclear whether this scales, as complex bugs could have many blame commits, potentially overwhelming the agent with noisy signals.
This paper presents HAFixAgent, a history-aware agent for APR that integrates history heuristics (inspired by HAFix (Shi et al., 2025)) to enable the agent to leverage historical context alongside bash command tools during its autonomous execution. To evaluate the effectiveness of HAFixAgent, we apply it to all 854 bugs in Defects4J (Just et al., 2014) (Java) and all 501 bugs in BugsInPy (Widyasari et al., 2020) (Python), two widely used benchmarks for evaluating APR techniques. In this study, we address the following research questions (RQs):
  • RQ0. To What Extent Is Historical Information Available in Real-World Bugs? Before integrating repository history, we first establish if relevant historical information is broadly available and scalable for use in APR. We analyze both Defects4J and BugsInPy to measure the availability and concentration of historical information. Our findings confirm that 71.1% (Defects4J) and 87.4% (BugsInPy) of bugs have available historical information. In Defects4J, 99.5% of blameable bugs map to a single blame commit, while BugsInPy shows more scattered blame (58.4% single-commit).
  • RQ1. How Effective Is HAFixAgent at Fixing Real-World Bugs? We empirically compare HAFixAgent against two state-of-the-art baselines replicated on the same LLM (DeepSeek-V3.2-Exp) for a fair comparison, and perform a controlled ablation (history vs. non-history) on both Defects4J and BugsInPy. HAFixAgent outperforms RepairAgent by +56.6% and BIRCH-feedback by +47.1%. Historical context further improves repair by +4.4% on Defects4J and +38.6% on BugsInPy, with the three history heuristics solving complementary sets of bugs.
  • RQ2. How Efficient and Robust Is History-Aware Repair? A practical concern is whether history-aware repair remains effective under imperfect fault localization and whether it inflates agent cost. We evaluate robustness by applying line-shift noise (+1/+3/+5) to fault localization on a stratified sample of 100 Defects4J bugs, and analyze cost/efficiency across both benchmarks. We find that history context provides increasing resilience under noisy FL, and does not significantly inflate cost.
The main contributions of our paper are as follows:
  • The design and implementation of HAFixAgent, a novel history-aware agentic approach for APR that enriches its decision-making loop with deep repository context.
  • A practical and automated method for constructing historical context, featuring three distinct blame-derived historical heuristics and a robust fallback strategy for any agentic APR system.
  • A large-scale empirical evaluation on 854 Defects4J (Java) and 501 BugsInPy (Python) bugs, with same-LLM baseline replication, providing strong evidence that history (1) is broadly available across languages, (2) significantly improves agent effectiveness, (3) provides increasing resilience under noisy fault localization, and (4) does not inflate cost.
Paper organization.
Section 2 provides background information. Section 3 presents the preliminary study on history availability and distribution. Section 4 describes our architecture design of HAFixAgent. Section 5 presents our case study design. Section 6 presents the results. Section 7 discusses the implications. Section 8 outlines threats to the validity of our study. Section 9 reviews related work. Section 10 concludes this paper.
2. Background
In this section, we introduce the background information for our study.
2.1. Mining Repository History for APR
A persistent challenge in LLM-based APR is providing the model with relevant, high-quality context (Prenner and Robbes, 2024; Sintaha et al., 2023). While many systems focus on the current (buggy) code snapshot, the repository’s version control history offers a rich, longitudinal source of context. Decades of Mining Software Repositories (MSR) research have established that historical data is a powerful asset. Hassan (2006) demonstrated the broad utility of mining repositories to assist developers. More specifically, Sliwerski et al. (2005) showed that analyzing the change-inducing commit can help locate faults. Adams et al. (2010) used history to identify co-evolving files or crosscutting concerns, providing a map of code dependencies not visible in a static snapshot.
Building on these MSR insights, researchers have long attempted to create APR systems that directly leverage history. For example, Le et al. (2016) pioneered this by creating a repair model based on patterns mined from previous, successful human-generated patches. This idea of using bug-to-fix patterns became a staple in pre-LLM repair techniques. Similarly, Kamei et al. (2013) reviewed the state of using historical data for the related task of fixing build failures, confirming the value of this approach.
Even in the modern LLM era, this principle of leveraging history-as-experience is re-emerging to enhance agentic repair. For example, EXPEREPAIR (Mu et al., 2025) designs a dual-memory system that augments an LLM’s context by leveraging historical repair experiences from previously resolved issues from the same repository. Likewise, SWE-Exp (Chen et al., 2025) proposes an experience-driven framework where the agent learns from the past issue-resolution trajectories (i.e., history) to improve its current issue repair.
These agent-based approaches show the value of history but rely on similar bugs or issues occurring before in the code repository or past trajectories. A more direct and practical heuristic is available through tools like git blame. For any given file, git blame annotates each line with the specific commit that last modified it. The commit identified by this command serves as a powerful heuristic. This choice aligns with SZZ’s core assumption: given a bug-fixing change, SZZ blames the lines modified or deleted by the fix in the pre-fix revision and treats the last change to those lines as a likely bug-introducing commit (Sliwerski et al., 2005; Kim et al., 2006). Analyzing the patch (diff) and commit message from this single, relevant change can provide critical clues about the code’s recent evolution and developer intent. Our prior work, HAFix (Shi et al., 2025), validated this exact approach, demonstrating that injecting this blame-derived context significantly improves repair for single-line bugs. This paper builds directly on that finding: we aim to explore the hypothesis that integrating this same focused historical context into a modern agentic workflow (Section 2.2) can scale these benefits to handle more complex, multi-location repairs.
2.2. Agentic Software Engineering
The emergence of powerful large language models (LLMs), such as GPT-5 (OpenAI, 2025a) and DeepSeek-V3 (DeepSeek-AI et al., 2024), has catalyzed a paradigm shift in software engineering, moving from static, one-shot code generation to dynamic, autonomous agent-based systems. An LLM-based agent is a system that couples the core reasoning and planning capabilities of an LLM with essential components like memory (to maintain state) and a set of tools (to interact with an environment) (Wang et al., 2024). Unlike a simple prompt-and-response model, an agent operates within an iterative execution loop, often conceptualized as a Reason-Act (ReAct) cycle (Yao et al., 2023). This loop allows the agent to formulate a plan, execute an action, observe the outcome, and then autonomously react to that feedback to progress toward a complex, high-level goal.
This iterative, stateful, and interactive approach is the foundation of Agentic Software Engineering (ASE), an emerging field that marks a paradigm shift for software engineering in the era of foundation models (Hassan et al., 2024). The vision of ASE is to create autonomous ”AI Teammates” or ”AI Software Developers” that can operate as generalist partners, capable of handling complex, end-to-end software development tasks (Hassan et al., 2025; Li et al., 2025b). The software engineering domain is a particularly fertile ground for agents because developer tasks are goal-oriented (e.g., ”fix this bug,” ”implement this feature”) and require interaction with a rich, well-defined tool ecosystem, including file systems, build tools, compilers, test runners, and version control systems (Roychoudhury, 2025).
Within the broad field of ASE, a primary and highly active research area has been software debugging and repair, specifically for resolving bugs and GitHub issues. This has led to a new generation of agentic APR systems. For example, SWE-agent (Yang et al., 2024) introduced an Agent-Computer Interface (ACI) that equips an agent with simple, developer-centric tools (such as file navigation, editing, and searching) to autonomously resolve real-world GitHub issues. Similarly, the OpenHands (Wang et al., 2025) seeks to build generalist agents that can execute complex tasks by reasoning over a plan and invoking tools. In the specific APR domain, RepairAgent (Bouzenia et al., 2025) was a pioneering work that provided an LLM with a dedicated toolset to autonomously read and extract code, search code, generate patches, and react to test feedback. Around the same time, systems like AutoCodeRover (Zhang et al., 2024c) further enhanced agents by integrating context from program analysis, such as API calls, to improve repair performance.
These systems represent the state-of-the-art in agent-based program repair. However, a common limitation is that the context these agents leverage is typically limited to the current, static snapshot of the codebase and the immediate feedback from the test suite. They generally lack awareness of the codebase’s historical and evolutionary context: how and why files have changed over time. HAFixAgent, the APR system proposed in this paper, is designed to bridge this exact gap by enriching this agentic loop with actionable, historical context derived from version control history.
3. Preliminary Study (RQ0): To What Extent Is Historical Information Available in Real-World Bugs?
In this section, we provide the motivation, approach, and results for the preliminary study.
3.1. Motivation
Providing LLM agents with informative context is essential for effective repair (Prenner and Robbes, 2024). While most context engineering approaches for automated program repair (APR) focus on knowledge obtained from a given snapshot of a code base (Prenner and Robbes, 2024; Haque et al., 2025; Fakhoury et al., 2024; Chen et al., 2024; Xia et al., 2023b; Li et al., 2025a), only recently APR techniques were proposed that leverage code history (Shi et al., 2025; Ehsani et al., 2025). Inspired by the domain of mining software repositories (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013), prior work on single-line APR (Shi et al., 2025) leverages context related to the last commit that touched the bug location, as the latter commit commonly is assumed to be the bug-introducing commit and hence has valuable knowledge about the introduced bug. Finding this commit for single-line bugs is straightforward (as obtained using the git blame command), as they only comprise one buggy line.
Unfortunately, more complex bugs touching one hunk, multiple hunks in one file or multiple hunks across files potentially would require combining blame commit information of many different lines, potentially drowning the latter information within an LLM’s token window, and hence reducing the performance of historical context for APR. In such a case, an APR technique would need to aggregate and rank multiple blame history sources, increasing prompt size, latency, and the risk of distraction. Furthermore, for a given buggy snapshot of the code base, blame commits do not exist for all buggy lines. For instance, when fault localization identifies the fault location as an insertion point (where code is missing) rather than an existing line, git blame cannot be applied, as it only annotates existing code. Notably, this process focuses solely on the buggy lines identified by fault localization, which does not depend on the actual bug fix.
This preliminary RQ studies how much of a problem the scale of blame history is for single-hunk, multi-hunk and multi-file bugs: (i) how often does a blame commit exist for buggy lines? and (ii) across how many unique commits is the historical blame information of these types of bugs spread? The results will provide clear empirical insights about the effort required to adapt history-based context to more complex types of bugs.
3.2. Approach
3.2.1. Dataset
To study the availability and distribution of historical information in real-world bugs across languages, we choose two established benchmarks. Defects4J (v3.0.1) (Just et al., 2014) is a standard and widely-used benchmark for evaluating recent APR techniques (Xia and Zhang, 2022; Jiang et al., 2023; Hossain et al., 2024; Lutellier et al., 2020; Jiang et al., 2021; Ye and Monperrus, 2024; Bouzenia et al., 2025). It contains 854 Java bugs alongside test cases for each bug collected from 17 different open-source projects, spanning various domains such as data visualization (Chart), compiler construction (Closure), and date/time utilities (Time). BugsInPy (Widyasari et al., 2020) is a benchmark of 501 Python bugs collected from 17 popular open-source projects, including data processing (pandas), web frameworks (fastapi, tornado), and utility libraries (tqdm, thefuck). BugsInPy complements Defects4J by covering a dynamically typed language with different development practices and commit patterns.
Consistent with common practices in the APR field (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we use the developer-written patch provided by each benchmark to identify the buggy lines (i.e., the lines in the buggy snapshot that were modified or deleted in the fix). Importantly, the approach does not rely on the actual bug fixes themselves, but solely on the locations of the buggy lines. These identified lines are the only lines on which git blame is subsequently computed. We then run git blame on these lines to obtain the most recent commits modifying those lines. Finally, we de-duplicate the commit hashes obtained for each bug, count each bug’s unique blame commits, and compute the distribution by bug category.
3.2.2. Bug Category Terminology
To categorize bugs based on the structure of their bug fix commits, we first define a hunk as a single, contiguous block of edits (i.e., additions, deletions, or modifications) applied to a specific region in the source code, similar to earlier work (Nashid et al., 2025). A patch file can contain one or more such hunks. Furthermore, following prior work on bug categorization (Nashid et al., 2025; Bouzenia et al., 2025; Jiang et al., 2023; Xia and Zhang, 2024), we classify all 854 bugs into four mutually exclusive categories based on the number and location of these hunks:
  • Single-Line (SL): The bug fix commit edits exactly one line, contained within a single hunk in a single file.
  • Single-Hunk (SH): The bug fix commit edits one contiguous hunk in one file, and consists of two or more line changes.
  • Single-File-Multi-Hunk (SFMH): The bug fix commit contains two or more distinct hunks, but all hunks are confined to a single file.
  • Multi-File-Multi-Hunk (MFMH): The bug fix commit contains hunks that span two or more different files.
We also categorize bugs by blame availability:
  • Blameable: At least one edited line in the bug fix commit is a deletion or modification relative to the buggy code snapshot, so git blame can return the most recent commit touching such lines.
  • Blameless: The bug fix commit only inserts new lines or files, with no deletions or modifications, so no line maps to a prior commit and hence git blame is not applicable.
Unlike prior work that either nests SL within SH (Jiang et al., 2023) or collapses SFMH and MFMH into one multi-hunk class (Nashid et al., 2025), we keep all four categories separate for finer granularity.
Figure 1. Distribution of unique blame commits per bug category on Defects4J (a) and BugsInPy (b). Defects4J blame is almost entirely concentrated in a single commit, while BugsInPy shows substantial multi-commit blame for complex bug categories.
3.3. Results
Historical blame commit information is broadly available, even for complex bugs. The majority of bugs in both datasets are blameable, i.e., have at least one blame commit across all buggy lines: 71.1% in Defects4J and 87.4% in BugsInPy. In Defects4J, complex multi-hunk bugs (SFMH and MFMH), which account for 64.8% of the dataset, maintain high history availability: 69.2% of SFMH and 80.5% of MFMH bugs are blameable. SH has the lowest rate at 52.2%, as single-hunk bugs more frequently are pure insertions. In BugsInPy, blame availability is even higher across all categories, with SL at 92.7%, SFMH at 91.6%, and MFMH at 92.3%, while SH remains the lowest at 73.8%, consistent with the Defects4J pattern. These results indicate that historical blame data is commonly available across bug categories and languages, supporting history-aware APR techniques. For bugs with no blame commit (28.9% in Defects4J, 12.6% in BugsInPy), a workaround would be required.
Blame concentration differs sharply across languages: Defects4J is usually single-commit, whereas BugsInPy is often multi-commit. Figure 1(a) shows that 70.7% (604 out of 854) of Defects4J bugs map to exactly one unique blame commit, and only 0.4% (3 out of 854) have two or more. Manual inspection confirmed that multi-hunk bugs typically trace back to a single common commit, as their buggy lines largely overlap in history. In contrast, BugsInPy blame is more scattered, with 51.1% (256 out of 501) of bugs mapping to a single blame commit and 36.3% (182 out of 501) having two or more. Multi-commit proportion rises with bug complexity: 0% for SL, 14.3% for SH, 49.0% for SFMH, and 71.4% for MFMH. Among blameable bugs, only 58.4% in BugsInPy are single-commit, compared to 99.5% in Defects4J. This contrast likely reflects differences in development practices: Python projects in BugsInPy tend to have finer-grained commits touching individual lines across sessions, while Defects4J’s Java projects more often introduce related changes in atomic commits.
Figure 2. HAFixAgent architecture and workflow.
4. HAFixAgent
Based on the findings of the preliminary study, we propose a new APR approach called HAFixAgent that exhibits the following three design decisions:
  1. (1)
    a minimal agent-based architecture designed to iteratively handle complex, multi-step reasoning and actions (e.g., code exploration, patch validation) required for bugs that go beyond simple, single-line fixes.
  2. (2)
    the agent has access to a git blame tool, inspired by the findings in Section 3, and leverages the blame commit as the primary source of historical context. When multiple blame commits exist, an LLM-based judge selects the most relevant one.
  3. (3)
    the agent includes a fallback mechanism to obtain the relevant history for blameless bugs where no buggy line has an associated blame commit.
Unlike agent-based APR systems such as RepairAgent (Bouzenia et al., 2025), which employ pre-defined Python tool APIs, HAFixAgent is intentionally minimal and history-aware. It treats repository history from previous snapshots as an important source of context that augments the current code snapshot, and it operates within a compact, guarded loop over a small set of standard bash tools (e.g., grep, sed, find). This design choice is critical: it minimizes ambiguity from complex and high-level tool interactions, allowing us to more directly attribute performance gains to the quality of the historical context rather than to sophisticated tool exploration. Additionally, the design is consistent with recent commercial coding assistants such as Claude Code (Anthropic, 2025) and OpenAI Codex (OpenAI, 2025b).
Figure 2 overviews the architecture and its three primary modules. The workflow proceeds as follows:
  • Context Builder: This module prepares the full context for the agent. It assembles the Metadata Context 1 and uses its History Extractor to retrieve 2 and format 3 the Historical Context. All context is then organized into the User Prompt and System Prompt 4.
  • Agent Execution Loop: This module controls the iterative repair. It starts with the LLM query 5, which generates an action 6. This action is sent to the Tools module 7. The loop then receives the raw feedback 8 from the tools, tracks the test result 9, and checks termination guards 10 to decide whether to continue. If so, the feedback is routed to the next LLM query $1$, repeating the cycle.
  • Tools: This module provides an isolated sandbox environment for action execution. It receives an action (e.g., search, edit, test) from the execution loop 7, then returns the resulting feedback (e.g., stdout/stderr) to the loop 8.
4.1. Context Builder
4.1.1. Metadata Context
The user prompt contains three items: the human-written bug report, the initial set of failing tests, and the fault localization context, which specifies the buggy files and lines. These metadata are used only as non-history context and are rendered in step 1.
4.1.2. History Extractor
The History Extractor module (step 2 in Figure 2) finds a relevant blame commit for every bug. We use the buggy lines identified through the fault localization metadata to obtain the blame commit, since the historical context in that blame commit may reveal the root causes of the bug (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013). The module handles blameable and blameless bugs differently:
  • Blameable bugs. For bugs involving modified or deleted lines, we execute git blame on all identified lines and collect the union of all unique commit hashes. As demonstrated in our preliminary study (Section 3), this union is often a single commit (99.5% of blameable Defects4J bugs), though BugsInPy bugs more frequently have multiple blame commits (41.6%). When multiple commits exist, we use an LLM-as-a-judge to select the most relevant one (detailed in Section 5.1).
  • Blameless bugs (fallback strategy). To enable these blameless bugs (i.e., add-only fixes) to also benefit from historical context, we design a fallback strategy. Specifically, we employ a fallback strategy to blame the nearest executable code line (e.g., an executable line rather than a comment or symbol) within the five lines preceding the insertion. This strategy is based on the assumption that new code is often added near existing, related logic, making the most recent commit to that local area a relevant source of historical context.
4.1.3. Historical Context
Once the blame commits are determined by the History Extractor, step 3 then mines relevant historical context from the identified blame commits by leveraging one of the three top-performing historical heuristics from the HAFix approach (Shi et al., 2025) (which focused on SL bugs only):
  • All functions’ names from the blame commit (fn_all): includes the names of all functions from all co-changed files in the blame commit. This setting captures a broad structural view of the co-evolved functions within the same history commit.
  • Historical function before and after (fn_pair): which contains the before and after snapshot of the function code body that includes the buggy line. This historical context helps the agent understand how the buggy function evolved and provides possible clues about the root cause of the bug.
  • File-level diff (fl_diff): includes the diff patch obtained from the git diff command in the blame commit. This configuration exposes the exact code modification across files, allowing the agent to reason about fine-grained edits that may have caused the bug.
In step 4, HAFixAgent integrates the history context, which was extracted in the previous step 3 using one of the three heuristics from the blamed commit. The history is injected directly into the User Prompt (as shown in Figure 2), with truncation safeguards (e.g., truncating an exceptionally large file diff) applied to ensure that the total prompt fits within the LLM’s context window. This process leads to our four context configurations: the agent operates either with the non-history context only (from Section 4.1.1) or with the non-history context augmented by one of the three history heuristics (fn_all, fn_pair, or fl_diff).
4.2. Agent Execution Loop
HAFixAgent’s execution loop follows an observe-act-feedback paradigm, inspired by frameworks like ReAct (Yao et al., 2023) and those used in recent software engineering agents (Yang et al., 2024). The loop proceeds with the following steps: 5 the LLM receives the current state (including the full conversation history of prior actions and observations); 6 the agent parses the model output to extract a tool invocation (i.e., a bash command, which the system prompt constrains the model always to provide) as the next action; 7 the command runs in a containerized tool sandbox that exposes repair utilities and dataset tools (e.g., search, edit and compile); 8 stdout and stderr are collected from the output of command execution as feedback; 9 compilation and test outcomes are captured and parsed; 10 the loop terminates on success or when a termination guard (e.g., step limit, cost limit, or timeout) is reached; otherwise, $1$ the observation is appended to the conversation history and fed back to the LLM for the next reasoning step.
4.3. Tools
Table 1. Bash tools used by HAFixAgent.
Tool Description Usage Example
Repair Operations
grep Search for patterns in files; Quick file overview of classes, interfaces, and method signatures grep -n "class|interface|public.*(" file.java
sed -n Read specific line ranges from files; extract targeted context around fault locations progressively sed -n ’45,65p’ BuggyClass.java (reads lines 45-65)
sed -i In-place file editing; make precise, targeted fixes to source code sed -i ’s/oldCode/newCode/’ file.java
find Locate files in the repository by name or pattern find . -name "*Test.java"
head / tail Reconstruct files for complex multi-line edits; extract beginning/end of files for rebuilding head -n 50 file.java > temp && cat fix.txt >> temp
Project Commands
compile Compile the project to verify setup and check for compilation errors after edits compile
test Run relevant/failing tests; verify bug fix effectiveness test -r
General Commands
echo Signal task completion when all tests pass echo COMPLETE_REPAIR_SIGNAL
&& / || Chain multiple commands together; execute sequentially or conditionally sed -n ’10,20p’ A.java && sed -n ’30,40p’ B.java
Table 1 lists the bash tools that HAFixAgent may invoke during the execution loop. The agent emits exactly one bash command per step, with optional short chaining via && or ||. These actions are executed between steps 7 and 8 in the loop, and the same toolset is embedded in the system prompt (Appendix A).
HAFixAgent intentionally exposes a simple but powerful set of bash tools together with project commands like compile and test. A key design choice is that the historical context is not an interactive tool, but is instead modularized and prepared by the History Extractor and injected directly into the agent’s context before the execution loop begins. This decouples history retrieval from agent actions, which allows the agent’s toolset to remain simple, modular, scalable and focused on repair-related operations (e.g., read, edit and test).
The tools in Table 1 fall into three groups. This specific set of tools was chosen to be minimal yet complete, providing the fundamental capabilities (file operations, project compilation, and testing) required for the iterative repair process while relying on universal, standard Linux utilities. (i) File operations: grep gives a quick structural overview, sed -n reads targeted windows around fault locations with progressive expansion, sed -i applies precise edits, head/tail supports multiline reconstruction by splicing, and find locates candidate files. (ii) Project commands: compile command checks build health after edits; test command runs only the relevant tests to provide fast feedback in the execution loop. (iii) General commands: a completion sentinel (echo COMPLETE_REPAIR_SIGNAL) signals success, and simple chaining helps combine a small read or edit with an immediate compile or test.
In practice, the loop alternates between these actions: inspect and read around the localized lines, apply a minimal edit, compile, run relevant tests, and iterate until all tests pass. Termination occurs in two ways: (1) the agent, after running a successful test command, self-terminates by emitting the echo COMPLETE_REPAIR_SIGNAL, or (2) the agent execution loop detects that a guard (step, cost, or timeout) has been reached and halts the process.
5. Case Study Design
In this section, we describe the case study design for evaluating HAFixAgent.
5.1. Prompt Context Preparation
Consistent with the preliminary study, we employ the entire 854 bugs of Defects4J to prepare the data needed for evaluation.
Metadata collection.
To prepare the non-history context as defined in Section 4.1.1, we need bug metadata such as bug reports, failing tests, and fault locations. For Defects4J, we mine bug reports from issue tracker links provided by the benchmark. For 18 cases from the Chart project without a clear bug issue link, we use the fixing commit message as the bug description, and to prevent potential data leakage, we manually verified these commit messages to ensure they do not contain explicit fix instructions. For BugsInPy, we intentionally omit bug descriptions to avoid data leakage from commit messages, relying on fault locations and failing tests only. Failing tests are collected from each benchmark’s test framework. Consistent with common practices in the field of program repair (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we assume perfect fault localization and provide fault localization context, which is derived from the developer’s bug fix (following the process in Section 3.2), directly into the agent’s context.
Historical data collection.
Motivated by the preliminary study in Section 3, which shows that blame commits are broadly available in both datasets, we collect history per bug as follows:
  • Blameable with one unique commit: Run git blame on deleted or modified lines to obtain the blame commit, as defined in Section 4.1.3.
  • Blameable with multiple commits: In Defects4J, only three bugs fall in this case; in BugsInPy, 182 bugs have two or more blame commits. We use the LLM-as-a-judge (the same LLM as HAFixAgent) to select the single most relevant commit. The model is prompted with the set of blameable lines and asked to choose the line which is most likely related to the bug’s root cause.
  • Blameless with zero commits: Apply the nearest line fallback strategy from Section 4.1.3 to select a blame commit.
Given the selected blame commit, we extract and construct data of fn_all, fn_pair, and fl_diff as defined in Section 4.1.3.
5.2. Baselines
We compare HAFixAgent with two recent APR techniques as baselines, i.e., RepairAgent (Bouzenia et al., 2025) and BIRCH-feedback (Nashid et al., 2025). RepairAgent is an autonomous LLM-based agent equipped with pre-defined API tooling. The original study used GPT-3.5-turbo and evaluated on 835 bugs (from Defects4J v1.2 and v2), reporting 186 plausible and 164 correct fixes. We replicate RepairAgent on DeepSeek-V3.2-Exp and evaluate on all 854 Defects4J bugs, using the plausible metric (test-passing patches) for a direct, same-metric comparison. BIRCH-feedback targets multi-hunk repairs. The original study used o4-mini and reported results on 371 multi-hunk bugs. We replicate their best-performing configuration with feedback on DeepSeek-V3.2-Exp and evaluate on the same 371 overlapping multi-hunk bugs (244 SFMH + 127 MFMH).
To eliminate the model confound and ensure a fair comparison, we replicate both baselines on the same LLM used by HAFixAgent (DeepSeek-V3.2-Exp), rather than relying on their originally reported numbers obtained with different models.
5.3. Metrics
In line with previous studies (Chen et al., 2021a; Parasaram et al., 2025; Ehsani et al., 2025; Nashid et al., 2025; Silva and Monperrus, 2025), we report Plausible@1 to evaluate the correctness based on whether the generated patch passes all test cases. Formally:
(1) $\mathrm{Plausible@1}=\frac{1}{|\mathcal{B}|}\sum_{b\in\mathcal{B}}\mathrm{TestPass}\!\left(\text{patch}^{\,b}_{1}\right)$
where $\mathcal{B}$ denotes the set of bugs and $\mathrm{TestPass}(\cdot)\in\{0,1\}$ returns $1$ if the patch compiles and all tests pass, and 0 otherwise. We generate one patch per bug using standard decoding settings, with the temperature set to 0.0 for enhancing reproducibility.
Complementing Plausible@1, which measures the pass rate as a proportion of all bugs, we also report #Pass (the absolute count of bugs passing the test suite), similar to prior work (Nashid et al., 2025). To quantify the unique contribution of diverse historical context, we additionally report #Unique Pass, to show the number of bugs solved by a configuration using historical context but not by its non-history counterpart. We visualize overlaps using Venn diagrams to illustrate multi-way complementarity.
For efficiency, we report the average number of agent steps per bug, the average cost per bug in USD computed from token usage, and the cost stratified by outcome (successful versus failed fixes). To assess the statistical significance of differences in cost and number of steps, we compare the distributions across the four context configurations. Since the data is paired (same bugs, different configurations) and not assumed to be normally distributed, we employ the Friedman test (Friedman, 1937), a non-parametric test for multiple comparisons. If the Friedman test is significant ($p<0.05$), we conduct pairwise Wilcoxon signed-rank tests (Wilcoxon, 1992) for post-hoc analysis to compare each history-aware configuration against the non-history configuration. To control for multiple comparisons (three pairs), we apply the Bonferroni correction (Dunn, 1961) and set the significance threshold $\alpha$.
5.4. Implementation
HAFixAgent is implemented in Python and builds on mini-swe-agent (Yang et al., 2024) for the execution loop, which aligns with our lightweight design. While mini-swe-agent (and swe-agent) is designed naturally to support solving PR issue benchmarks such as SWE-Bench (Jimenez et al., 2024), it lacks the capability of historical context management (e.g., history localizing and retrieval), the repairing strategies and tools specifically for resolving Defects4J bugs. It is an open-source, lightweight framework with about 100 lines of core loop code, enabling seamless integration of historical context extraction and bash tools orchestration during runtime. All experiments (HAFixAgent, RepairAgent, and BIRCH-feedback) use DeepSeek-V3.2-Exp (DeepSeek-AI et al., 2024; DeepSeek-AI, 2025) via OpenRouter with temperature 0.0. Prompts are rendered at runtime with jinja2.
We run HAFixAgent on Ubuntu 20.04, with each bug executing in an isolated Docker container. For Defects4J, the sandbox is built from the released Defects4J image, with the compile and test tools (Section 4.3) wrapping defects4j compile and defects4j test -r. For BugsInPy, we build pre-warmed Docker images per project with the correct Python environment, and the test tool wraps bugsinpy-test. For each bug, we cap the agent’s loop at 50 steps, enforce a per-bug cost guard of \$1 USD, and set a 1-hour timeout. After each run, the container is automatically cleaned, while all generated patches and logs are persisted for analysis.
6. Results
In this section, we provide the motivation, approach, and results for each of our research questions.
6.1. RQ1: How Effective Is HAFixAgent at Fixing Real-World Bugs?
6.1.1. Motivation
Prior work (Shi et al., 2025) shows that injecting historical context improves LLM repair for single-line bugs under simple prompting. History carries heuristic signals that are hard to infer from a static snapshot: the fault introducing change, coevolving functions and files, and often intent hints in commit diffs and messages. RQ0’s finding that blame commits are broadly available suggests that historical context could scale to more complex bug types. However, RQ0 also reveals that blame concentration differs across benchmarks: while Defects4J blame is highly concentrated (99.5% single-commit), BugsInPy shows more scattered blame (41.6% multi-commit), raising the question of whether a single-commit strategy generalizes across languages.
This RQ rigorously validates the benefit of blame-derived context on SH, SFMH and MFMH bugs, in the context of our proposed HAFixAgent, across both Java and Python benchmarks. To thoroughly validate our approach, we first establish HAFixAgent’s practical viability by benchmarking it against state-of-the-art (SOTA) baselines on the same LLM to determine if our history-aware method is competitive. Second, we isolate the specific impact of our core hypothesis by evaluating the agent across different context configurations (i.e., with versus without history) on both Defects4J and BugsInPy. This analysis is essential to quantify the precise contribution of historical context, understand how gains vary by bug category, and determine whether the effect generalizes across languages. This motivates the following questions: (a) is HAFixAgent competitive against SOTA baselines on the same LLM? (b) what is the performance impact of adding historical context to an APR agent, and (c) how do these gains vary by bug category and across benchmarks?
6.1.2. Approach
To ensure a fair comparison, we replicate the two SOTA baselines on the same LLM used by HAFixAgent (DeepSeek-V3.2-Exp), eliminating the model confound present in prior work. We compare HAFixAgent against RepairAgent (Bouzenia et al., 2025) on all 854 Defects4J bugs and BIRCH-feedback (Nashid et al., 2025) on 371 multi-hunk bugs (244 SFMH + 127 MFMH), matching their original scope. To analyze performance across different bug types, we categorize results by our four bug types (SL, SH, SFMH, MFMH). We pay special attention to the multi-hunk categories (SFMH and MFMH), as their need for coordinated changes may make them more sensitive to historical context.
In addition, we evaluate the four context configurations defined in Section 4.1 across all 854 Defects4J bugs and all 501 BugsInPy bugs. The four configurations are HAFixAgent-non-history, HAFixAgent-fn_all, HAFixAgent-fn_pair, and HAFixAgent-fl_diff. Blameless bugs (28.9% in Defects4J, 12.6% in BugsInPy) use the fallback strategy (Section 4.1.3), ensuring every bug has meaningful historical context.
All experiments use the same LLM, identical parameters, and a consistent containerized runtime. We report both Plausible@1 (computed via termination checks in Section 4.2) and #Pass (the total number of bugs with patches passing all tests). To quantify the specific benefit of historical context, we also report #Unique Pass for the number of bugs repaired by a history-aware configuration that were not repaired by the non-history configuration.
6.1.3. Results
Table 2. Same-LLM effectiveness comparison on Defects4J (all methods use DeepSeek-V3.2-Exp). Numbers show plausible patches passing the full test suite. Bold = best per category.
(a) All 854 Defects4J bugs: HAFixAgent vs RepairAgent
Cat. Total HAFixAgent RepairAgent
non-history fn_all fn_pair fl_diff
SL 167 137 136 136 140 121
SH 134 97 96 108 103 68
SFMH 425 238 251 249 254 141
MFMH 128 50 47 43 48 18
Total 854 522 530 536 545 348
(b) 371 multi-hunk bugs: HAFixAgent vs BIRCH-feedback
Cat. Common HAFixAgent BIRCH-fb
non-history fn_all fn_pair fl_diff
SFMH 244 123 125 133 126 85
MFMH 127 49 46 42 47 34
Total 371 172 171 175 173 119
HAFixAgent substantially outperforms both RepairAgent and BIRCH-feedback on the same LLM. On all 854 Defects4J bugs (Table 2(a)), the best HAFixAgent configuration (fl_diff) repairs 545 bugs versus 348 for RepairAgent, a +56.6% improvement. Even HAFixAgent-non-history (522), which uses no historical context, outperforms RepairAgent by +50.0%, demonstrating that the agent architecture itself is already more effective. The gap is largest for complex bugs: on SFMH, HAFixAgent-fl_diff repairs 254 versus 141, and on MFMH, HAFixAgent-non-history repairs 50 versus 18.
On the 371 multi-hunk bugs evaluated by BIRCH-feedback (Table 2(b)), the best HAFixAgent configuration (fn_pair) repairs 175 bugs versus 119 for BIRCH-feedback, a +47.1% improvement. By category, HAFixAgent-fn_pair repairs 133 SFMH and 42 MFMH bugs, compared with 85 and 34 for BIRCH-feedback. Since all methods now use the same LLM, these improvements directly reflect differences in repair strategy rather than model capability.
Table 3. History ablation on Defects4J (854 Java bugs) and BugsInPy (501 Python bugs). Each cell shows #Pass (Plausible@1). Superscript = #Unique bugs fixed by that config but not by non-history. Bold = best per category.
Dataset Category Total non-history fn_all fn_pair fl_diff
Defects4J SL 167 137 (82.0%) 136 (81.4%) +10 136 (81.4%) +10 140 (83.8%)+11
SH 134 97 (72.4%) 96 (71.6%) +16 108 (80.6%)+19 103 (76.9%) +18
SFMH 425 238 (56.0%) 251 (59.1%) +67 249 (58.6%) +69 254 (59.8%)+68
MFMH 128 50 (39.1%) 47 (36.7%) +13 43 (33.6%) +13 48 (37.5%) +12
Total 854 522 (61.1%) 530 (62.1%) +106 536 (62.8%) +111 545 (63.8%)+109
BugsInPy SL 82 74 (90.2%) 73 (89.0%) +4 73 (89.0%) +7 71 (86.6%) +5
SH 126 93 (73.8%) 100 (79.4%) +20 105 (83.3%)+20 93 (73.8%) +16
SFMH 202 54 (26.7%) 148 (73.3%) +107 137 (67.8%) +95 157 (77.7%)+118
MFMH 91 51 (56.0%) 56 (61.5%) +23 59 (64.8%)+17 53 (58.2%) +21
Total 501 272 (54.3%) 377 (75.2%)+154 374 (74.7%) +139 374 (74.7%) +160
Figure 3. Venn diagrams comparing bug repair overlap across 4 context configurations on Defects4J (top) and BugsInPy (bottom). Each number shows the count of bugs uniquely fixed by that intersection of configurations.
Table 4. Median cost (USD) and agent steps for successful repairs, with Friedman test p-values for cost on the matched set (N = bugs fixed by all 4 configs). Friedman tests for steps were non-significant on all categories (all $p\geq 0.05$). Bold p-values indicate significance ($p<0.05$); marks configs with significantly higher cost than non-history (pairwise Wilcoxon, Bonferroni $\alpha=0.0167$).
Median Cost (USD) Median Steps
Dataset Cat. N non-history fn_all fn_pair fl_diff non-history fn_all fn_pair fl_diff Fried. p
Defects4J SL 118 0.005 0.008 0.005 0.008 12 14 13 16 <<0.001
SH 72 0.016 0.019 0.015 0.014 23 25 24 22 0.184
SFMH 140 0.016 0.020 0.016 0.021 26 28 26 30 0.122
MFMH 22 0.029 0.026 0.023 0.023 34 29 32 30 0.142
BugsInPy SL 59 0.027 0.029 0.020 0.033 19 19 18 18 0.214
SH 61 0.030 0.031 0.034 0.036 23 18 23 21 0.200
SFMH 27 0.070 0.037 0.055 0.055 26 24 24 23 0.157
MFMH 20 0.062 0.064 0.066 0.060 29 30 31 30 0.204
On Defects4J, history improves repair by +4.4% (545 vs 522), with strong complementarity across categories. Table 3 (top) shows the Defects4J ablation. Across all 854 bugs, the best history configuration (fl_diff) repairs 545 versus 522 for non-history (+4.4%). By category, fl_diff performs best on SL (140 fixes, 83.8%) and SFMH (254 fixes, 59.8%), while fn_pair leads on SH with 108 fixes (80.6%). MFMH reveals a trade-off: non-history achieves the highest total (50), yet history configurations collectively add 24 unique fixes not found by non-history, outweighing the 8 unique non-history-only fixes (Figure 3). This indicates that history provides an orthogonal signal even when it does not improve the aggregate count.
On BugsInPy, history improves repair by +38.6% (377 vs 272), with SFMH gaining +190.7% (54 to 157). Table 3 (bottom) shows the BugsInPy ablation. The best history configuration (fn_all) repairs 377 out of 501 bugs (75.2%) versus 272 for non-history (54.3%), a gain of +105 bugs (+38.6%). The SFMH category drives the largest improvement: fl_diff repairs 157 SFMH bugs (77.7%) compared with only 54 for non-history (26.7%), a +190.7% improvement. History also helps on SH (+12 bugs with fn_pair) and MFMH (+8 bugs with fn_pair). The substantially larger history benefit on BugsInPy compared to Defects4J may reflect differences in benchmark characteristics, such as language, project complexity, or the lower baseline performance of the non-history configuration on BugsInPy (54.3% vs 61.1%), leaving more room for history to contribute.
The three history configurations are complementary, solving distinct sets of bugs. The Venn diagrams in Figure 3 show large shared cores in each category, yet each heuristic contributes nontrivial unique fixes. On Defects4J SFMH, fn_all, fn_pair, and fl_diff each add 21, 19, and 19 unique cases beyond the 140-bug shared core. On BugsInPy SFMH, the shared core is only 27 bugs, with history configurations expanding the fixable set dramatically (72 bugs in the three-way history overlap alone). This pattern confirms that fn_all (function-name co-evolution), fn_pair (before/after semantics), and fl_diff (fine-grained edits) capture complementary signals. As shown in the summary of Figure 3, the union of all three history configurations fixes 684 Defects4J bugs versus 522 for non-history (+31.0%), and 463 BugsInPy bugs versus 272 (+70.2%), with the largest per-category gains on SFMH and MFMH.
6.2. RQ2: Efficiency and Robustness
6.2.1. Motivation
While RQ1 shows that historical context improves repair effectiveness, two practical concerns remain. First, cost is a primary concern for agent systems, as iterative tool use can accumulate substantial token and time budgets, even on simple bugs (Bouzenia et al., 2025; Xia and Zhang, 2024). Recent evaluations also highlight the risk of expensive failures where unresolved attempts consume several times more tokens than successful ones (Fan et al., 2025). Since HAFixAgent integrates additional contextual information into each prompt, it is important to verify that the performance gains from RQ1 do not incur significant cost overhead. Second, all RQ1 experiments assume perfect fault localization (FL), which is unrealistic in practice (Yang et al., 2024; Fan et al., 2023). In practice, FL tools produce imprecise locations, and when blame is computed on incorrect lines, irrelevant commits may be retrieved, potentially negating the history benefit. This RQ addresses two concerns: (a) whether adding historical context inflates cost, and (b) whether the history benefit persists under noisy FL.
6.2.2. Approach
We evaluate inference cost (USD) and agent reasoning steps across all four configurations on Defects4J (854 bugs) and BugsInPy (501 bugs). Inference cost is computed from token usage converted to USD using API pricing rates for DeepSeek-V3.2-Exp. A run terminates upon a successful repair or by hitting one of the predefined limits (50 steps, \$1 USD cost, or 1-hour timeout), as defined in Section 5.4. To ensure fair comparison, we report cost and step metrics on the subset of bugs fixed by all four configurations, and assess statistical significance using Friedman tests followed by pairwise Wilcoxon signed-rank tests with Bonferroni correction.
To evaluate robustness under imperfect FL, we simulate noise by applying positive line shifts of +1, +3, and +5 to the fault locations of a stratified sample of 100 Defects4J bugs (25 per category: SL/SH/SFMH/MFMH, randomly sampled with seed=42, excluding the Chart project). The shift affects the entire pipeline: the agent receives shifted locations in its prompt, blame is computed on the shifted lines (potentially retrieving different or irrelevant commits), and history context is extracted from the shifted blame. The existing perfect-FL results from RQ1 serve as the shift=0 baseline. We run all 4 configurations at each shift level, yielding $100\times 4\times 3=1{,}200$ additional runs. Running the full datasets across all shifts would require 16,260 runs and is prohibitively expensive, so we adopt stratified sampling with equal representation across bug categories to maintain coverage while keeping costs manageable.
6.2.3. Results
Figure 4. Success rate degradation under noisy fault localization (line shifts +1, +3, +5) on 100 Defects4J bugs (25 per category). Shift=0 uses perfect FL from RQ1.
Successful repairs converge well before the step cap. Table 4 reports median agent steps and cost by configuration and category on both datasets. On Defects4J, the median steps for successful repairs is 12 to 16 in SL, 22 to 25 in SH, 26 to 30 in SFMH, and 29 to 34 in MFMH, all well below the 50-step cap. BugsInPy shows a similar pattern (18 to 19 in SL, 18 to 23 in SH, 23 to 26 in SFMH, 29 to 31 in MFMH). Harder categories require more steps but still finish well under the cap on both benchmarks.
History does not significantly increase step counts or cost. To rigorously assess whether historical heuristics significantly alter repair costs compared to non-history, we conducted Friedman tests on the matched set of bugs fixed by all four configurations (Table 4, Friedman p column). On Defects4J, only SL shows a significant cost difference ($p<0.001$); pairwise Wilcoxon tests with Bonferroni correction ($\alpha=0.0167$) show that fn_all ($p=0.0022$) and fl_diff ($p=0.001$) cost slightly more than non-history on SL. For SH, SFMH, and MFMH, costs are statistically comparable (all $p\geq 0.05$). Notably, the median costs for successful MFMH repairs are comparable or lower with history (e.g., Defects4J: non-history \$0.029 vs fn_pair \$0.023; BugsInPy: non-history \$0.062 vs fl_diff \$0.060).
The same cost pattern holds on BugsInPy. Table 4 reports median cost and steps for both datasets, along with Friedman test p-values. On Defects4J, only SL shows a significant cost difference ($p<0.001$); for SH, SFMH, and MFMH, costs are statistically comparable across configurations. On BugsInPy, no category shows significant differences (all $p\geq 0.15$), confirming that history does not inflate cost on either benchmark. Notably, on BugsInPy SFMH, the non-history configuration is the most expensive (\$0.070 median) while history configurations cost \$0.037 to \$0.055, suggesting that history helps the agent converge faster on complex bugs.
Combining history configurations increases success but also increases cost, with diminishing returns. On Defects4J, two-config combinations such as fn_pair +fl_diff reach 90.4% on SL at \$0.041, close to the four-config union (95.2% at \$0.086). On BugsInPy, the pattern is similar but with higher absolute costs. Across both datasets, two or three configurations give the best success per dollar, while adding the fourth yields small extra gains at substantially higher cost.
Under noisy FL, history provides increasing resilience: on SFMH, non-history collapses to 0% at shift=+5 while history maintains 40 to 56%. Figure 4 shows success rates as FL noise increases. The most pronounced effect appears on SFMH: the non-history configuration drops from 64% at shift=0 to 0% at shift=+5. In contrast, history configurations maintain 40 to 56% success at shift=+5 (fl_diff and fn_all both at 56%), indicating that historical context acts as a safety net when FL is imprecise. The history advantage on SFMH grows with noise, from 4-4% at shift=0 to +56+56% at shift=+5.
On SH, a similar pattern appears at moderate noise: at shift=+3, the best history configurations (fn_all and fn_pair, both 88%) outperform non-history (68%) by 20%. At shift=+5, all configurations degrade, but history still leads (36% vs 24%). SL bugs remain robust across configurations, consistent with their simplicity, while MFMH shows no clear trend due to the small sample size and inherent difficulty. Although individual data points may fluctuate with 25 bugs per category, the overall degradation trend and the widening gap between history and non-history remain consistent. These results show that history-aware repair not only benefits from precise fault locations but also provides contextual value that becomes increasingly important as FL precision decreases.
7. Discussion
Performance Regressions. We observe 32 cases where a history configuration underperforms the non-history one or misses non-history fixes (Figure 3). For instance, in MFMH, the non-history configuration attains 50 while fn_all, fn_pair, and fl_diff reach 47, 43, and 48, respectively, even though each contributes 12 to 13 unique fixes that the non-history does not solve. This pattern aligns with LLM evidence that more context can hurt: models are easily distracted by irrelevant additions (Shi et al., 2023), and performance often declines when relevant facts sit in the middle or deeper parts of long prompts (Levy et al., 2024). In our setting, extra historical snippets can dilute salient context and reduce Plausible@1 on some bugs that don’t need historical context.
Heuristics Complementarity. A key finding from RQ1 (Figure 3) is that the three historical heuristics are complementary, not redundant, with each repairing a distinct set of bugs. This suggests that different bug types are susceptible to different forms of historical context: fl_diff excels at fine-grained edit patterns (SL/SH), fn_pair provides semantic context showing function evolution (complex logic errors), and fn_all captures co-evolutionary patterns across functions (multi-hunk bugs). The strongest benefits appear for SFMH bugs, which involve multiple related changes within a single file. Because the blame commit typically captures the entire change context for such bugs (as shown by the single-commit concentration in RQ0), all three heuristics provide highly relevant information about the bug-introducing pattern. This complementarity also explains the practical trade-off identified in RQ2: combining two or three heuristics yields the most cost-effective performance.
Plausible vs. Correct Patches. Our evaluation uses plausible patches (passing the full test suite) as the success metric, consistent with prior work (Nashid et al., 2025; Ehsani et al., 2025; Silva and Monperrus, 2025). Manual verification of over 3.5k plausible patches (4 configurations $\times$ 2 datasets) is prohibitively expensive at this scale. Our evaluation focuses on relative improvement rather than absolute correctness: the same plausible metric and same LLM are used consistently across all comparisons, including both replicated baselines and the history vs. non-history ablation, ensuring fair comparison. More broadly, as LLM and agent-based APR systems generate plausible patches at unprecedented scale, designing automated patch correctness validation becomes a pressing open problem for the community.
8. Threats to Validity
Internal Validity. DeepSeek-V3.2-Exp is available as open weights, but the pretraining corpus is not disclosed to our knowledge. As a result, we cannot verify if the specific projects or bugs in Defects4J are included in the pretraining data. However, our evaluation focuses on the relative effectiveness improvement of different historical heuristics, rather than their absolute performance. For instance, if HAFixAgent-fl_diff fixes a bug that HAFixAgent-non-history does not fix, this improvement is attributed to the benefit of historical context, regardless of whether the LLM has seen the bug before. Furthermore, the substantially larger history benefit on BugsInPy (+38.6%) compared to Defects4J (+4.4%) supports this interpretation: if data leakage explained the gains, both benchmarks should show similar history benefits, since the same LLM and same ablation design are used. The larger BugsInPy effect suggests that history provides genuine signal that is more impactful when the LLM has less memorized knowledge about the benchmark.
Our main evaluation assumes perfect fault localization, a common setup in LLM-based APR studies (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025). To mitigate this threat, RQ2a evaluates robustness under noisy FL (line shifts of +1/+3/+5), showing that history-aware configurations maintain effectiveness even when FL is imprecise. However, our noise model (uniform positive shifts) does not capture all real-world FL errors, such as entirely wrong file predictions.
External Validity Our study evaluates on Defects4J (854 Java bugs) and BugsInPy (501 Python bugs). While this covers two languages and 34 projects, results may not transfer to other ecosystems where project size, test adequacy, and repository practices differ. The history profiles already differ between our two benchmarks (e.g., 99.5% single-commit in Defects4J vs. 58.4% in BugsInPy), suggesting that history-aware techniques may need adaptation per ecosystem. We expose a dataset-agnostic interface behind the history extractor and release all code and scripts for replication. Evaluating on repository-level benchmarks such as SWE-bench (Jimenez et al., 2024) is future work.
HAFixAgent is instantiated with one agent loop, a narrow bash tool set, specific prompts, and a single LLM with fixed step, cost, and time guards. Different agents, models, or decoding policies may change absolute numbers; API-level determinism does not guarantee identical runs. To reduce this threat, we use DeepSeek-V3.2-Exp and replicate all baselines on the same LLM.
9. Related Work
In this section, we discuss related work about the traditional automated program repair, LLM and agent based automated program repair, and in-context learning for automated program repair.
9.1. Traditional Automated Program Repair
Traditional APR formulates patch generation as a search or constraint problem over program transformations guided by tests or specifications. Search-based systems evolve patches with genetic or heuristic search over mutation operators and validate candidates against the test suite, exemplified by GenProg and successors (Le Goues et al., 2012; Le et al., 2016; Wen et al., 2018). Semantics or constraint-based repair derives patches by solving synthesis constraints obtained from symbolic execution or program analysis, as in SemFix (Nguyen et al., 2013) and Angelix (Mechtaev et al., 2016), and with targeted condition synthesis such as Nopol (Xuan et al., 2017). Template-based repair uses human-designed or mined fix schema that match buggy contexts and instantiate edits such as TBar (Liu et al., 2019b) and AVATAR (Liu et al., 2019a). A complementary line reuses in-project code fragments as ingredients, such as ssFix (Xin and Reiss, 2017) and CAPGEN (Wen et al., 2018). These approaches established core pipelines for localization, candidate generation, and test-based validation, but their coverage is constrained by operator or pattern design and by search scalability.
Learning-based APR before LLMs frames repair as code-to-code translation or edit prediction trained on a large amount of code corpus. Early neural systems include DeepFix (Gupta et al., 2017) for syntax errors and sequence-to-sequence repair, such as SequenceR (Chen et al., 2021b), with later improvements from model ensembling and syntax guidance in CoCoNuT (Lutellier et al., 2020), CURE (Jiang et al., 2021), and Recoder (Zhu et al., 2021). Template mining and ranking were also brought to production, for example, Facebook’s Getafix (Bader et al., 2019) for static-analyzer warnings and the end-to-end SapFix (Marginean et al., 2019) pipeline that proposed developer-reviewed patches. Together, traditional APR demonstrates effective repair across bug classes while revealing recurring limitations in computation cost, low accuracy and overfitting to tests. These limitations have motivated LLM and agent-based APR that learn fixes from data and plan tool-driven repair loops, which we introduce next.
9.2. LLM and Agent based Automated Program Repair
Large language models (LLMs) have significantly advanced Automated Program Repair (APR), moving from direct prompting to sophisticated, multi-step agentic systems. Early evidence shows code LLMs already outperform prior learning-based APR and benefit from fine-tuning and careful evaluation on standard benchmarks (Jiang et al., 2023; Xia et al., 2023b). This led to a variety of prompting and interaction strategies. For example, conversation-driven repair simulates a developer’s debugging process by interleaving patch generation with test feedback, achieving a strong balance of cost and accuracy (Xia and Zhang, 2024). Other approaches focus on improving the model’s reasoning, such as self-directed repair, which uses a chain-of-thought process to gather knowledge before fixing (Yin et al., 2024). The plastic surgery line revisits ingredient reuse by aligning LLMs with project-specific code (Xia et al., 2023a). Beyond pure prompting, hybrid and template-guided methods constrain or scaffold the patch space with analysis or repair templates to improve plausibility and correctness (Li et al., 2025a; Huang et al., 2025).
Agent-based APR systems represent a further step, shifting from single-shot patch generation to autonomous, tool-using workflows. Pioneering general-purpose software agents, like SWE-agent (Yang et al., 2024) and OpenHands (Wang et al., 2025), established a powerful paradigm by equipping LLMs with tools to interact with a repository, such as editors, shell commands, and test runners. Building on this, repair-focused agents have integrated more specialized tools. For instance, some employ search and fault localization to narrow down the buggy code (Bouzenia et al., 2025; Zhang et al., 2024c), while others explore multi-agent collaboration to divide the debugging task (Lee et al., 2024), with empirical studies beginning to map this rapidly evolving design space (Meng et al., 2024). A recent trend also focuses on improving agent performance on complex, repository-level tasks by incorporating memory and experience, allowing agents to learn from prior repair trajectories (Mu et al., 2025; Chen et al., 2025). These lines generally rely on local code and runtime outputs, with limited study of commit history as a first-class context inside the loop, which is the focus of our work.
9.3. In-context Learning for Automated Program Repair
A core question for LLM-based APR is which context to provide and how to structure it. Prior work shows that adding bug-related facts such as error messages, stack traces, and failing tests improves repair, and that the choice of facts matters (Parasaram et al., 2025; Xia and Zhang, 2024). Studies on local context indicate that models are sensitive to how much and which code surrounds the edit region (Prenner and Robbes, 2024). Other work leverages natural language bug reports or transforms descriptions into edits, and explores execution traces to guide patching (Fakhoury et al., 2024; Haque et al., 2025). Repository-level evaluations report that larger, realistic tasks remain challenging and require broader context curation (Chen et al., 2024).
Beyond local signals, historical context is receiving attention. HAFix (Shi et al., 2025) demonstrates that blame-derived commit data is an effective, lightweight historical signal for fixing more bugs, motivating history-aware agent designs at Defects4J scale. Layered knowledge injection systematically adds project and repository knowledge to prompts and improves bug fixing on broader-context scenarios (Ehsani et al., 2025). Built on top of HAFix, HAFixAgent advances this direction by bringing blame-driven history into the agentic workflow and quantifying its effect on single-hunk and multi-hunk bugs.
10. Conclusion and Future Work
This paper presented HAFixAgent, a history-aware agentic approach that injects blame-derived repository context into an observe-act-verify loop to address the more complex single- and multi-hunk bug categories. A preliminary study on 854 Defects4J (Java) and 501 BugsInPy (Python) bugs shows that blame commit history is widely available (71.1% and 87.4% blameable, respectively), with Defects4J blame highly concentrated in a single commit (99.5%) while BugsInPy shows more scattered blame (58.4% single-commit).
Using the same LLM for all experiments including replicated baselines, HAFixAgent outperforms RepairAgent (+56.6%) and BIRCH-feedback (+47.1%) on Defects4J. Historical context further improves repair by +4.4% on Defects4J and +38.6% on BugsInPy, with the largest gains on SFMH bugs (+190.7% on BugsInPy). Under noisy fault localization, history provides increasing resilience, particularly on SFMH where non-history collapses to 0% while history maintains up to 56%. History does not significantly inflate agent cost on either benchmark, and the three complementary heuristics each solve distinct bug subsets.
Two promising directions for future work emerge from our findings. First, HAFixAgent currently uses an LLM judge to select a single blame commit; for BugsInPy bugs where 41.6% have multiple blame commits, combining context from several commits could further improve repair effectiveness. Second, evolving HAFixAgent into a multi-agent architecture (Rafi et al., 2024; Lin et al., 2025; Lee et al., 2024) with specialized History, Fixer, and Reviewer agents could improve repair of complex MFMH bugs where a single agent’s context window is easily overwhelmed.
Appendix A HAFixAgent Prompt
System Prompt of HAFixAgent. The values inside placeholder will be rendered in runtime.